### v3

# thanks: corresponding author

Abstract
While quantum state tomography (QST) remains the gold standard for benchmarking and verifying quantum devices, it requires an exponentially large number of measurements and classical computational resources for generic quantum many-body systems, making it impractical even for intermediate-size quantum devices. Fortunately, many physical quantum states often exhibit certain low-dimensional structures that enable the development of efficient QST. A notable example is the class of states represented by matrix product operators (MPOs) with a finite matrix/bond dimension, which include most physical states in one dimension and where the number of independent parameters describing the states only grows linearly with the number of qubits. Whether a sample efficient quantum state tomography protocol, where the number of required state copies scales only linearly as the number of parameters describing the state, exists for a generic MPO state still remains an important open question.
In this paper, we answer this fundamental question affirmatively by using a class of informationally complete positive operator-valued measures (IC-POVMs)—including symmetric IC-POVMs (SIC-POVMs) and spherical $t$-designs— focusing on sample complexity while not accounting for the implementation complexity of the measurement settings. For SIC-POVMs and (approximate) spherical 2-designs, we show that the number of state copies to guarantee bounded recovery error of an MPO state with a constrained least-squares estimator depends on the probability distribution of the MPO under the POVM but scales only linearly with $n$ when the distribution is approximately uniform. For spherical $t$-designs with $t\geq 3$, we prove that only a number of state copies proportional to the number of independent parameters in the MPO is sufficient for a guaranteed recovery of any state represented by an MPO. Moreover, we propose a projected gradient descent (PGD) algorithm to solve the constrained least-squares problem and show that it can efficiently find an estimate with bounded recovery error when appropriately initialized.
1 Introduction
The state of a quantum system composed of $n$ qudits (which are $d$-level quantum systems; qubits have $d=2$) is fully described by a density matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$ with unit trace and is positive semidefinite (PSD) [1]. To reconstruct or estimate the density matrix, quantum measurements need to be performed on numerous identical copies of the state. Any such physical measurement is characterized by a Positive Operator-Valued Measure (POVM), which can be represented by a set of PSD matrices $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ that collectively sum to the identity matrix, i.e., $\sum_{k=1}^{K}\bm{A}_{k}={\bf I}_{d^{n}}$. Each matrix $\bm{A}_{k}$ ($k=1,\dots,K$) in the POVM corresponds to a potential measurement outcome, and the probability of obtaining that outcome is given by $p_{k}=\operatorname{trace}(\bm{A}_{k}\bm{\rho})$. The probabilistic nature of quantum measurements necessitates multiple measurements (say $M$) with the same POVM to obtain statistically accurate estimates $\widehat{p}_{k}=\frac{f_{k}}{M}$ of each $p_{k}$, where $f_{k}$ denotes the number of times the $k$-th outcome is observed in the $M$ measurements. Disregarding statistical errors, $\{p_{k}\}$ can be considered as $K$ linear measurements of the state $\bm{\rho}$. From a machine learning perspective, we can refer to $\{p_{k}\}$ and their empirical estimates $\{\widehat{p}_{k}\}$ as population and empirical measurements of the state, respectively. The goal of quantum state tomography (QST) [2] is to find the density matrix from empirical measurements with high accuracy. As the density matrix provides full information about the quantum state, QST remains the gold standard for benchmarking and verifying quantum devices.
Extensive research efforts, both in algorithmic development [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and theoretical analysis [15, 16, 17, 18, 19, 20], have been devoted to QST. For generic quantum states, the required state copies must grow exponentially with the number of qudits. This sample complexity can be improved for low-rank density matrices [21, 19, 17], but it still grows exponentially with $n$. To implement QST effectively on current quantum computers with over 100 qubits [22, 23, 24], it is important to exploit more compact structures within practical states of interest. A notable example is the class of states represented by matrix product operators (MPOs) [^1] MPOs representing physical quantum states are also termed matrix product density operators (MPDOs), specifically designed for representing mixed states via density matrices. In contrast, matrix product states (MPSs) are used to represent pure states. . The density matrix elements of such states can be expressed as the following matrix product [25, 26]
$\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},$ (1)
where $i_{k}$ ($k=1,2,\cdots n$) labels a basis of the $k$-th qubit, and $\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}}$ with $r_{0}=r_{n}=1$. The dimensions $\bm{r}=(r_{1},\ldots,r_{n-1})$ are often called the bond dimensions  [^2] It is also common to simply call $\overline{r}=\max\{r_{1},\ldots,r_{n-1}\}$ the bond dimension. of the MPO in quantum physics, though we may also call them the MPO ranks. An MPO comprises $nd^{2}$ matrices, each with a dimension of at most $\overline{r}\times\overline{r}$ where the matrix dimension $\overline{r}=\max_{\ell}r_{\ell}$ determines the efficiency of the representation. Thus, an MPO can be efficiently represented with $O(nd^{2}\overline{r}^{2})$ number of parameters. The decomposition (1) is also mathematically equivalent to the tensor train (TT) decomposition used for compact representation of large tensors [27].
Examples of states that can be represented by MPOs include ground states of most quantum systems in one dimension [28], general Hamiltonians with decaying long-range interactions [29], and those generated by noisy quantum computers, where the noise could also limit the amount of quantum entanglement and thus enable an efficient state representation [26]. Several approaches, such as maximum likelihood estimation [30, 31] and tensor train cross approximation [32], have empirically demonstrated the scalable number of state copies when employing MPO in QST. Nevertheless, the theoretical analysis of this advantage has remained ambiguous in previous studies. Recently, the work [33] successfully established a theoretical guarantee on the recovery error of an MPO state using multiple Haar random projective measurements. To be specific, for any given $\epsilon>0$, [33] proves that with $O(nd^{2}\overline{r}^{2}\log n)$ different sets of Haar random projective measurement bases and $O(n^{3}d^{2}\overline{r}^{2}\log n/\epsilon^{2})$ total state copies, a properly constrained least-squares minimization using the empirical measurements stably can recover the ground-truth state with $\epsilon$-closeness in the Frobenius norm with high probability. While this result formally shows that MPOs can be reconstructed using polynomial number of state copies with regard to $n$, it is not sample efficient as the scaling $n^{3}$ is much larger than the degrees of freedom ($O(n)$) in the MPOs. In addition, this work [33] provides a projected gradient descent (i.e., iterative hard thresholding) algorithm for performing the QST in practice, but does not provide a formal guarantee for that algorithm.
1.1 Main contributions
Our work aims to address two fundamental questions regarding the QST of states represented by MPOs:
  1. 1.
    Is there a measurement setting that can provably recover an MPO with a number of state copies proportional to the number of parameters in the MPO?
  2. 2.
    If so, can we develop a provably correct algorithm to reconstruct MPOs from the obtained measurements?
Table 1: Comparison with existing work for estimating $n$-qudits MPOs (with bond dimension $\overline{r}$) to achieve $\epsilon$-closeness in the Frobenius norm. The result for SIC/2-designs depends on $\gamma(\bm{\rho})=K\cdot\max_{k}p_{k}$, i.e., the probability distribution of measuring the state $\bm{\rho}$ with the corresponding POVM.
POVM ##\## POVMs ##\## total state copies
Haar random ([33, Theorem 5]) $\Omega(nd^{2}\overline{r}^{2}\log n)$ $O(\frac{n^{3}d^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
SIC/$2$-designs (Theorem 1)             $1$ $O(\frac{nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n}{\epsilon^{2}})$
$t$-designs, $t\geq 3$ (Theorem 2)             $1$ $O(\frac{nd^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
We provide positive answers to both the above questions by utilizing informationally complete POVMs (IC-POVM). An IC-POVM is defined by the property that every quantum state is uniquely determined by its measurement statistics. Beyond QST, IC-POVMs with special ties find applications in quantum cryptography [34], quantum fingerprinting [35], and are relevant to foundational studies of quantum mechanics [36, 37]. It eliminates the inherent randomness in Haar random POVMs that are often used to establish efficient sample complexity [21, 19]. Without this inherent randomness, we can more easily analyze the effects of statistical errors in the recovery error bound and the design provable optimization algorithms.
We begin by studying recovery guarantees for MPOs from empirical quantum measurements (physical measurements containing statistical errors) obtained using a class of IC-PVOMs including the rank-one symmetric IC-POVM (SIC-POVM) [38] and more broadly spherical $t$-design POVMs that are induced by (approximate) spherical $t$-designs [39, 40]. The first contribution of this paper—presented in Section 3—is a new sample complexity bound (refers to the number of state copies) that matches the number of parameters in MPOs to guarantee bounded recovery error for a particular estimator—the solution to a constrained least-squares optimization problem. Specifically, for SIC-POVM and $2$-design POVMs, our result shows that $M=O(nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n/\epsilon^{2})$ number of state copies is sufficient to achieve $\epsilon$-accuracy in recovering the ground-truth state $\bm{\rho}$ in the Frobenius norm, where $\gamma(\bm{\rho})=K\cdot\max_{k=1}^{K}p_{k}$ captures the uniformity of the probabilistic distribution of the measurement. The term $\gamma(\bm{\rho})$ can be as small as 1, but in some cases, it may be as large as $d^{n}$; however, for many cases it is only polynomially large in $n$. The dependence of $\gamma(\bm{\rho})$ could be eliminated by using spherical $t$-designs with $t\geq 3$, for which we only need $M=O(nd^{2}\overline{r}^{2}\log n/\epsilon^{2})$ to ensure the sufficient condition. Excluding the logarithmic term, this bound exhibits linear dependence on $n$, $d^{2}$, and $\overline{r}^{2}$, aligning precisely with the degrees of freedom $O(nd^{2}\overline{r}^{2})$ in MPOs and is thus sample efficient. This result improves upon previous work [33] mentioned above using Haar random projective measurement bases that requires $\Omega(nd^{2}\overline{r}^{2}\log n)$ POVMs and $O(n^{3}d^{2}\overline{r}^{2}\log n/\epsilon^{2})$ total number of state copies. We summarize the main results and comparison with our previous result [33] in Table 1. Moreover, physical MPO states are often approximately low-rank, allowing the previous analysis to be extended to the trace distance. Specifically, to ensure a bounded recovery error in terms of trace distance for rank-$s$ MPO states, the required number of state copies remains proportional to the degrees of freedom, $O(nsd^{2}\overline{r}^{2})$, when using (approximate) spherical $t$-designs with $t\geq 2$.
The second contribution of this paper—presented in Section 4—is the development of a projected gradient descent (PGD) with guaranteed convergence analysis for solving the aforementioned constrained least-squares optimization problem. In each iteration, the algorithm performs a gradient update, followed by a projection back to the MPO format with a unit trace. Even without the constraint on the trace, computing the optimal projection onto the set of MPOs is challenging and could be NP-hard [41]. We employ a sequential approach called TT-SVD [27]—developed for computing TT decomposition which is equivalent to MPO—for computing an approximate projection. While TT-SVD is not optimal, it is computationally efficient and comes with a sub-optimal guarantee. Leveraging this guarantee, we show that, with an appropriate initialization, PGD converges linearly to a vicinity of the ground-truth state, with a distance matching the results in Table 1. In addition, we show that spectral initialization provides a valid starting point, ensuring the convergence of PGD. Our numerical experiments in Section 5 suggest that PGD can also efficiently find the underlying state even with a random initialization.
In this paper, we primarily focus on global spherical $t$-design POVMs with $t\geq 2$, as we theoretically demonstrate that stable recovery of constrained least-squares optimization for all MPO states can be achieved using a polynomial number of state copies. In practice, however, local measurements are widely employed in quantum state tomography due to their efficient implementation and lower computational complexity. Notably, when the Hilbert-Schmidt distance for local marginals is bounded, [42] has shown that the recovery error in terms of trace distance grows polynomially with $n$ for a translation-invariant state on an infinite one-dimensional chain. Nevertheless, the recent work in [43] reveals that recovering long-range entangled states–such as the general Greenberger-Horne-Zeilinger (GHZ) state, which admits an MPO representation–requires an exponential number of state copies when relying on local measurements. This finding highlights that, both theoretically and practically, local measurements are insufficient to guarantee the recovery of all MPO states using a polynomial number of state copies.
We note that our work is centered on the fundamental question of whether the number of state copies required for QST can be efficient (i.e. proportional to the intrinsic degrees of freedom) for recovering a structured quantum many-body state. We investigate this problem using IC-POVMs, focusing on sample complexity while not accounting for the implementation complexity of the measurement settings. In practice, the SIC-POVM or the spherical $t$-designs with $t\geq 2$ have no known efficient implementation using local quantum circuits. To partially address this practicality concern, we provide numerical evidence that local SIC-POVMs, which can be implemented efficiently and readily on current quantum hardware, can be used to perform QST for typical MPO states with $O(n)$ state copies using the proposed PGD algorithm in Section 5.
1.2 Notation
We use calligraphic letters (e.g., $\mathcal{X}$) to denote tensors, bold capital letters (e.g., $\bm{X}$) to denote matrices, bold lowercase letters (e.g., $\bm{x}$) to denote column vectors, and italic letters (e.g., $x$) to denote scalar quantities. Elements of matrices and tensors are denoted in parentheses. For example, $\mathcal{X}(i_{1},i_{2},i_{3})$ denotes the element in position $(i_{1},i_{2},i_{3})$ of the order-3 tensor $\mathcal{X}$. The calligraphic letter $\mathcal{A}$ is reserved for the linear measurement map. For a positive integer $K$, $[K]$ denotes the set $\{1,\dots,K\}$. The superscripts $(\cdot)^{\top}$ and $(\cdot)^{\dagger}$ denote the transpose and Hermitian transpose, respectively  [^3] As is conventional in the quantum physics literature (but not in information theory and signal processing), we use $(\cdot)^{\dagger}$ to denote the Hermitian transpose.. For two matrices $\bm{A},\bm{B}$ of the same size, $\left\langle\bm{A},\bm{B}\right\rangle=\operatorname{trace}(\bm{A}^{\dagger}\bm{B})$ denotes the inner product between them. $\|\bm{X}\|$ and $\|\bm{X}\|_{F}$ respectively represent the spectral norm and Frobenius norm of $\bm{X}$. $\sigma_{i}(\bm{X})$ is the $i$-th singular value of $\bm{X}$. For a vector $\bm{a}$ of size $N\times 1$, its $l_{n}$-norm is defined as $||\bm{a}||_{n}=(\sum_{m=1}^{N}|a_{m}|^{n})^{\frac{1}{n}}$. For two positive quantities $a,b\in\mathbb{R}$, the inequality $b\lesssim a$ or $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b\gtrsim a$ or $b=\Omega(a)$ represents $b\geq ca$ for some universal constant $c$. To simplify notations in the following sections, for an order-$n$ TT format with ranks $(r_{1},\dots,r_{n})$, we define $\overline{r}=\max_{i=1}^{n-1}r_{i}$.
2 Informationally Complete POVM
As described in the introduction, the population measurements $\{p_{k}\}$ of the density matrix $\bm{\rho}$ from a POVM $\{\bm{A}_{k}\}$ are linear measurements that can be described through a linear map $\mathcal{A}:\mathbb{C}^{d^{n}\times d^{n}}\rightarrow\mathbb{R}^{K}$ of the form
$\mathcal{A}(\bm{\rho})=\begin{bmatrix}\langle\bm{A}_{1},\bm{\rho}\rangle\\ \vdots\\ \langle\bm{A}_{K},\bm{\rho}\rangle\end{bmatrix}.$ (2)
If a POVM consists of at least $d^{2n}$ matrices such that we can form exactly $d^{2n}$ linearly independent matrices by linear combination, it is said to be an informationally complete POVM (IC-POVM) [44, 45, 46]. IC-POVMs consisting of exactly $d^{2n}$ elements are called minimal [47]. In this section, we introduce some common IC-POVMs, including symmetric IC-POVM and spherical $t$-designs, and present properties of the corresponding linear mappings $\mathcal{A}$.
Symmetric Informationally Complete POVM (SIC-POVM)
A particularly attractive and potentially useful IC-POVM is one that is symmetric, i.e., it has equal pairwise inner products between any pair of the POVM elements [38, 48]. A POVM that is informationally complete, minimal, and symmetric is called SIC-POVM.
Definition 1.
([49]) A Symmetric Informationally Complete POVM (SIC-POVM) consists of $d^{2n}$ rank-one POVM elements $\{\bm{A}_{k}\}$ that satisfies $(i)$ each with an equivalent trace $\operatorname{trace}(\bm{A}_{k})=\frac{1}{d^{n}}$ for all $k\in[d^{2n}]$, and $(ii)$ every pair of measurement operators satisfies $\langle\bm{A}_{k},\bm{A}_{j}\rangle=\begin{cases}\frac{1}{d^{2n}},&k=j,\\ \frac{1}{d^{2n}(d^{n}+1)},&k\neq j.\end{cases}$
The set of vectors comprising an SIC-POVM has been extensively studied in various contexts under different names [38], such as equiangular lines in linear algebra [50] and equalangular tight frames in information theory [51]. In quantum information, it has also found extensive applications, including quantum state tomography [52] and quantum cryptography [53]. SIC-POVMs have been numerically constructed for matrix dimensions ranging from $2$ to $151$ [54, 55], and for many higher dimensions, up to $39,604$ [56]. While it is widely believed that SIC-POVMs exist in all dimensions, no explicit construction of the SIC-POVM is known for an arbitrary dimension of the Hilbert space. The following result establishes stable embeddings of any Hermitian matrices from SIC-POVM in terms of $\|\mathcal{A}(\bm{\rho})\|_{2}^{2}$.
Lemma 1.
Suppose that $\{\bm{A}_{k}\}_{k=1}^{d^{2n}}$ form an SIC-POVM. Then for any Hermitian matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$, $\mathcal{A}(\bm{\rho})$ satisfies
$\|\mathcal{A}(\bm{\rho})\|_{2}^{2}=\sum_{k=1}^{d^{2n}}\langle\bm{A}_{k},\bm{\rho}\rangle^{2}=\frac{\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}(\bm{\rho}))^{2}}{d^{n}(d^{n}+1)}.$ (3)
The proof is provided in Appendix A. By using tools from compressive sensing literature [57, 58, 59, 60, 61], the previous work [33] provides a similar stable embedding result from Haar random projective measurement bases, but only for MPOs. In contrast, with POVM being informationally complete, (3) holds for any Hermitian matrices without relying on any randomness in the measurement operator. This allows us to develop tight sample complexity bound in Section 3 and iterative algorithms with guaranteed convergence in Section 4. As an immediate consequence of Lemma 1, for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, the difference $\bm{\rho}_{1}-\bm{\rho}_{2}$ remains Hermitian, and hence
$\|\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})\|_{2}^{2}$ $=$ $\frac{\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}+(\operatorname{trace}(\bm{\rho}_{1}-\bm{\rho}_{2}))^{2}}{d^{n}(d^{n}+1)}$ (4)
$=$ $\frac{\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}}{d^{n}(d^{n}+1)},$
where we use the property that $\operatorname{trace}(\bm{\rho}_{1})=\operatorname{trace}(\bm{\rho}_{2})=1$. Thus, SIC-POVM not only ensures distance measurements ($\mathcal{A}(\bm{\rho}_{1})\neq\mathcal{A}(\bm{\rho}_{2})$) as long as $\bm{\rho}_{1}\neq\bm{\rho}_{2}$, but also exactly preserves the distance of the two states in the measurement space, up to a universal scaling $1/d^{n}(d^{n}+1)$.
$t$-design POVMs
A more general set of IC-POVM are those induced by spherical $t$-designs.
Definition 2.
(Spherical $t$-designs [62, 15]). A finite set $\{\bm{w}_{k}\}_{k=1}^{K}\subset\mathbb{C}^{d^{n}}$ of normalized vectors is called a spherical quantum $t$-design if 444 $K\geq C_{d^{n}+\lfloor t/2\rfloor-1}^{\lfloor t/2\rfloor}\cdot C_{d^{n}+\lceil t/2\rceil-1}^{\lceil t/2\rceil}$ is necessary to form a spherical $t$-design[52, 63].
[…]
holds for any $s\leq t$, where the integral on the right hand side is taken with respect to the Haar measure on the complex unit sphere in $\mathbb{C}^{d^{n}}$.
[…]
Definition 3.
(Spherical $\delta$-approximate $t$-designs [67]). A finite set $\{\bm{w}_{k}\}_{k=1}^{K}\subset\mathbb{C}^{d^{n}}$ of normalized vectors is called a spherical quantum $\delta$-approximate $t$-design if
[…]
holds for any $s\geq 2$, where the integral on the right hand side is taken with respect to the Haar measure on the complex projective space. In addition, $\frac{1}{K}\sum_{k=1}^{K}\bm{w}_{k}\bm{w}_{k}^{\dagger}=\frac{1}{d^{n}}{\bf I}$.
The following result generalizes Lemma 1 to $t$-design POVMs and approximate $t$-design POVMs.
Lemma 2.
Suppose that $\{\bm{w}_{k}\}_{k=1}^{K}\subset\mathbb{C}^{d^{n}}$ forms an $\delta$-approximate $t$-design ($t\geq 2$). Let $\mathcal{A}$ be the linear map in (2) corresponding to the induced POVM $\{\bm{A}_{k}=\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger}\}$. Then for arbitrary Hermitian matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$, $\mathcal{A}(\bm{\rho})$ satisfies
$(1-\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}\leq\|\mathcal{A}(\bm{\rho})\|_{2}^{2}$
$\leq(1+\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}.$ (7)
The proof is given in Appendix B. Similar to (4), it follows from Lemma 2 that for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, $\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})$ satisfies
$\|\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})\|_{2}^{2}$ (8)
$\geq$ $(1-\delta)\frac{d^{n}(\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}_{1}-\bm{\rho}_{2}}))^{2})}{K(d^{n}+1)}$
$=$ $(1-\delta)\frac{d^{n}\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}}{K(d^{n}+1)}.$
3 Stable Recovery of Matrix Product Operators
Building upon the results that a (approximate) $t$-design POVM $\{\bm{A}_{1},\dots,\bm{A}_{K}\}\in\mathbb{C}^{d^{n}\times d^{n}}$ provides unique population measurements for any state, in this section, we investigate the stable recovery for MPOs of form (1) from empirical measurements acquired through such POVMs. Specifically, for an MPO $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ with MPO ranks $(r_{1},\dots,r_{n-1})$, we use the POVM to measure it $M$ times and take the average of the outcomes to generate empirical probabilities
$\widehat{p}_{k}=\frac{f_{k}}{M},\ k=1,\dots,K,$ (9)
where $f_{k}$ denotes the number of times the $k$-th output is observed. We denote ${\widehat{\bm{p}}}=\begin{bmatrix}\widehat{p}_{1}&\cdots&\widehat{p}_{K}\end{bmatrix}^{\top}$ as the empirical measurements obtained by the POVM, which are unbiased estimators of the population measurements $\bm{p}=\begin{bmatrix}p_{1}&\cdots&p_{K}\end{bmatrix}^{\top}=\mathcal{A}(\bm{\rho}^{\star})$, where $\mathcal{A}$ is the induced linear measurement operator formally defined in (2). To simplify the notation for the analysis, we denote by $\bm{\eta}=\begin{bmatrix}\eta_{1},\cdots,\eta_{K}\end{bmatrix}^{\top}=\widehat{\bm{p}}-\bm{p}$ the statistical error between the empirical probabilities and population probabilities.
With empirical measurements $\widehat{\bm{p}}$, we consider minimizing the following constrained least squares objective:
$\widehat{\bm{\rho}}=\operatorname*{\text{arg\leavevmode\nobreak\ min}}_{\bm{\rho}\in\mathbb{X}_{\bm{r}}}\|\mathcal{A}(\bm{\rho})-{\widehat{\bm{p}}}\|_{2}^{2},$ (10)
where we denote by $\mathbb{X}_{\bm{r}}$ the set of MPOs with MPO ranks $\bm{r}=(r_{1},\dots,r_{n-1})$:
$\begin{split}&\mathbb{X}_{\bm{r}}=\Big{\{}\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\ \bm{\rho}=\bm{\rho}^{\dagger},\operatorname{trace}(\bm{\rho})=1,\\ &\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},\\ &\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}},\ell=1,\dots,n-1,r_{0}=r_{n}=1\Big{\}}.\end{split}$ (11)
It is worth noting that one could consider imposing additional structure on the factors ${\bm{X}_{\ell}^{i_{\ell},j_{\ell}}}$ to ensure that $\bm{\rho}$ is PSD, as done in [68, eq. (3)]. However, the condition in [68, eq. (3)] is only sufficient, not necessary, for ensuring $\bm{\rho}$ is PSD and adding PSD constraint does not significantly reduce the number of degrees of freedom of elements in the set $\mathbb{X}_{\bm{r}}$. Hence, we focus on the set (11), which encompasses not only PSD matrices but also non-PSD matrices.
Assuming a global solution to (10) can be found, our objective is to study the performance of $\widehat{\bm{\rho}}$, and examine how the recovery error $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}$ scales with the size of the MPO (specifically in relation to the number of qudits $n$) and the number of measurements $M$. In order to achieve a stable estimation of the state with a polynomial number of measurements relative to $n$, we expect the recovery error exhibits polynomial growth rather than exponential growth with respect to $n$. Next section will provide iterative algorithm to solve (10).
Main results
We now provide a formal analysis of the recovery error $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}$. Since both SIC-POVM and exact $t$-design POVM can be viewed as specific instances of $\delta$-approximate $t$-design POVMs, the analysis will primarily focus on $\delta$-approximate $t$-design POVM. Using the fact that $\widehat{\bm{\rho}}$ is a global solution of (10) and the fact that $\bm{\rho}^{\star}\in\mathbb{X}_{\bar{r}}$, we have
00 0 $\leq$ $\|\mathcal{A}(\bm{\rho}^{\star})-\widehat{\bm{p}}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\widehat{\bm{p}}\|_{2}^{2}$ (12)
$=$ $\|\mathcal{A}(\bm{\rho}^{\star})-\mathcal{A}(\bm{\rho}^{\star})-\bm{\eta}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\mathcal{A}(\bm{\rho}^{\star})-\bm{\eta}\|_{2}^{2}$
$=$ $2\langle\mathcal{A}(\bm{\rho}^{\star})+\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle+\|\mathcal{A}(\bm{\rho}^{\star})\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})\|_{2}^{2}$
$=$ $2\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle-\|\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\|_{2}^{2},$
which further implies that
$\|\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\|_{2}^{2}\leq 2\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle.$ (13)
According to (8) for $\delta$-approximate $t$-design POVM, the left-hand side of the above equation can be further lower bounded by
$\|\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\|_{2}^{2}\geq(1-\delta)\frac{d^{n}\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}^{2}}{K(d^{n}+1)}.$ (14)
The challenge part is to deal with the right-hand side of (13). A straightforward application of the Cauchy–Schwarz inequality $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle\leq\|\bm{\eta}\|_{2}\cdot\|\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\|_{2}$ is insufficient for providing a tight result as $\|\bm{\eta}\|_{2}$ scales as $\frac{1}{\sqrt{M}}$ according to [33, eq. (35)]. Instead, we leverage the randomness of $\bm{\eta}$ and employ the following concentration bound for multinomial random variables.
Lemma 3.
([33, Lemma 3]) Suppose $(f_{1},\dots,f_{K})$ are mutually independent and follow the multinomial distribution $\operatorname{Multinomial}(M,\bm{p})$ where $\sum_{k=1}^{K}f_{i}=M$ and $\bm{p}=[p_{1},\cdots,p_{K}]$. Let $a_{1},\dots,a_{K}$ be fixed. Then, for any $t>0$,
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}a_{k}(\frac{f_{k}}{M}-p_{k})>t\right)$
$\leq e^{-\frac{Mt}{4a_{\max}}\min\bigg{\{}1,\frac{a_{\max}t}{4\sum_{k=1}^{K}a_{k}^{2}p_{k}}\bigg{\}}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}a_{k}^{2}p_{k}}},$
where $a_{\max}=\max_{k}|a_{k}|$.
We proceed by considering the following two cases, $2$-designs and $t$-designs with $t\geq 3$.
Case $(i)$: $2$-designs.
There are two potential technical difficulties in directly applying Lemma 3 to bound the term $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ by plugging in $a_{k}=\langle\bm{A}_{k},\widehat{\bm{\rho}}-\bm{\rho}^{\star}\rangle$: (1) $a_{k}$ in not independent to the random variables $\{f_{k}\}$ since $\widehat{\bm{\rho}}$ depends on the empirical measurements, and (2) we need to upper bound a term of form $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ originating from the exponential term in (3). We will address the first challenge by using a covering argument. For the second challenge, to utlize the upper bound for $\|\mathcal{A}(\bm{\rho}-\bm{\rho}^{\star})\|_{2}^{2}=\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}$ in Section 2, we capture the largest probabilities by
$\gamma(\bm{\rho}^{\star}):=K\cdot\max_{k}p_{k}=K\cdot\max_{k}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle.$ (16)
Since $\max_{k}p_{k}\geq\frac{1}{K}$, $\gamma(\bm{\rho}^{\star})\geq 1$ always holds. When the probability distribution is uniform, $\gamma(\bm{\rho}^{\star})$ can be small as $O(1)$; particularly, $\gamma(\bm{\rho}^{\star})=1$ when $p_{1}=\cdots=p_{K}=1/K$. On the other hand, if the probability distribution is spiky, $\gamma(\bm{\rho}^{\star})$ could approach $K$. We now summarize the main result as follows.
Theorem 1 (Stable recovery with $2$-designs).
Suppose $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate $2$-design POVMs. Consider an MPO state $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ with MPO ranks $(r_{1},\dots,r_{n-1})$. Then, with the empirical measurements $\widehat{\bm{p}}$ obtained by measuring the state $M$ times using the POVM and probability $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, the solution $\widehat{\bm{\rho}}$ of the constrained least squares (10) satisfies
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\lesssim\frac{d\overline{r}\sqrt{(1+\delta)n\gamma(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}}.$ (17)
The proof is given in Appendix C. Theorem 1 guarantees a stable recovery of the ground-truth state with a $\delta$-approximate $2$-design POVM when the number of state copies $M=O(nd^{2}\gamma(\bm{\rho}^{\star})\overline{r}^{2}\log n/\epsilon^{2})$. This condition highlights that the growth of $M$ is only polynomial in terms of the number of qudits $n$ when $\gamma(\bm{\rho}^{\star})$ is small. In particular, when $\gamma(\bm{\rho}^{\star})=O(1)$, we can achieve an efficient theoretical bound for $M$ that matches $O(nd^{2}\overline{r}^{2})$, the number of degrees of freedom of MPOs. On the other hand, when $\gamma(\bm{\rho}^{\star})$ is polynomial in terms of $n$, the order of $M$ still remains polynomial.
As mentioned above, $\gamma(\bm{\rho})\geq 1$ always holds. To further illustrate possible values for $\gamma(\bm{\rho})$, we consider the case of SIC-POVM. First consider an identity density matrix $\bm{\rho}=\frac{1}{d^{n}}{\bf I}_{d^{n}}=\frac{1}{d}{\bf I}_{d}\otimes\cdots\otimes\frac{1}{d}{\bf I}_{d}$ which can be written as an MPO with bond dimension $1$. For this case, $\langle\bm{A}_{k},\bm{\rho}\rangle=\frac{1}{d^{n}}\operatorname{trace}(\bm{A}_{k})=\frac{1}{d^{2n}}$ for all $k$, and hence $\gamma(\bm{\rho})=1$. On the other hand, when $\bm{\rho}=d^{n}\bm{A}_{k}$ for some $k$ (note that such a rank-one density matrix may not be an MPO), we have $\langle\bm{A}_{k},\bm{\rho}\rangle=\frac{1}{d^{n}}$ and $\langle\bm{A}_{j},\bm{\rho}\rangle=\frac{1}{d^{n}(d^{n}+1)}$ for all $j\neq k$, and hence $\gamma(\bm{\rho})=d^{n}$. Indeed, we can prove that $\gamma(\bm{\rho})\leq d^{n}$ always holds for any state $\bm{\rho}$ since $\langle\bm{A}_{k},\bm{\rho}\rangle\leq\|\bm{\rho}\|\operatorname{trace}(\bm{A}_{k})\leq\operatorname{trace}(\bm{A}_{k})=\frac{1}{d^{n}}$ for all $k$. However, we note that the above pathological case with spiky probability distribution may happen rarely in practice. For instance, consider a random density matrix $\bm{\rho}=\bm{u}{\bm{u}}^{\dagger}$, where [^5] Similar result also holds when $\bm{u}$ is randomly generated from the unit sphere according to the Haar measure. each entry in $\bm{u}\in\mathbb{C}^{d^{n}}$ in generated i.i.d. from a complex normal distribution $\mathcal{C}\mathcal{N}(0,\frac{1}{d^{n}})$. In this case, each term $\langle\bm{A}_{k},\bm{u}{\bm{u}}^{\dagger}\rangle$ has mean $\operatorname{\mathbb{E}}[\langle\bm{A}_{k},\bm{u}{\bm{u}}^{\dagger}\rangle]=\frac{\operatorname{trace}(\bm{A}_{k})}{d^{n}}=\frac{1}{d^{2n}}$. Moreover, since $\langle\bm{A}_{k},\bm{u}{\bm{u}}^{\dagger}\rangle$ is a second-order polynomial in the entries of Gaussian random vector, it is a subexponential random variable [69, Proposition 2.4] with subexponential norm  [^6] For a random variable $X$, its subgaussian norm and subexponential norm are respectively defined as $\|X\|_{\psi_{2}}=\inf\{t>0,\ \operatorname{\mathbb{E}}e^{\frac{|X|^{2}}{t^{2}}}\leq 2\}$ and $\|X\|_{\psi_{1}}=\inf\{t>0,\ \operatorname{\mathbb{E}}e^{\frac{|X|}{t}}\leq 2\}$. $\|\langle\bm{A}_{k},\bm{u}{\bm{u}}^{\dagger}\rangle-\frac{1}{d^{2n}}\|_{\psi_{1}}\leq\frac{c}{d^{n}}\|\sqrt{d^{n}}\bm{u}\|_{\psi_{2}}^{2}\|\bm{A}_{k}\|_{F}+\frac{1}{d^{2n}}\lesssim\frac{1}{d^{2n}}$ for some constant $c$. We can then invoke concentration inequality for subexponential random variable [70] to obtain that $\max_{k}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle\lesssim\frac{n}{d^{2n}}$ holds with probability at least $1-e^{-Cn}$ for some positive constant $C$. In other words, for such a random density matrix $\bm{\rho}$, with high probability the probability distribution under SIC-POVM is quite uniform as $\gamma(\bm{\rho})\lesssim n$.
We note that the dependence of $\gamma(\bm{\rho}^{\star})$ may arise from technical difficulties, and could be eliminated by more sophisticated analysis, which is left as future work. Instead, we exploit additional properties in the $3$-design to improve the bound.
Case $(ii)$: $t$-designs with $t\geq 3$
To provide a uniform bound for all MPOs without the dependence on $\gamma(\bm{\rho}^{\star})$, we can consider $3$-designs, utilizing the property that uniform sampling from a $3$-design mimics the first six moments of sampling uniformly according to the Haar measure. The following result establishes an improved guarantee for estimating MPO with $\delta$-approximate $t$-designs ($t\geq 3$).
Theorem 2 (Stable recovery with $t$-designs ($t\geq 3$)).
Suppose $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate $t$-design POVMs ($t\geq 3$) and $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ is an MPO state with MPO ranks $(r_{1},\dots,r_{n-1})$. If we measure the state $M$ times using the POVM, then with probability $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, the solution $\widehat{\bm{\rho}}$ of the constrained least squares (10) satisfies
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\lesssim\frac{d\overline{r}\sqrt{(1+\delta)n\log n}}{(1-\delta)\sqrt{M}}.$ (18)
The proof is given in Appendix D. With exact or approximate spherical $t$-design POVMs ($t\geq 3$), Equation 18 ensures a stable recovery of the ground-truth state when the number of state copies $M=O(nd^{2}\overline{r}^{2}\log n/\epsilon^{2})$. Notably, this uniform guarantee is applied to all MPOs. It is worth noting that the $n$-qubit Clifford group comprises a unitary (not spherical) 3-design POVM [71, 72, 73]. Sampling Clifford circuits uniformly at random reproduces the first $3$ moments of the full unitary group endowed with the Haar measure. Unfortunately, unitary 3-designs are not precisely equivalent to spherical 3-designs, thus we cannot efficiently construct spherical 3-designs. However, this work [74] demonstrates one method to generate spherical 3-designs by extracting them from unitary designs. On the other hand, the SIC-POVM is not a $3$-design POVM, hence the assumption of "uniformly" distributed probabilities may be indispensable.
While the solution $\widehat{\bm{\rho}}$ of (10) may be non-physical, we can impose additional PSD constraint to obtain a physical state without compromising the recovery guarantee in Theorem 1. Alternatively, we can simply project $\widehat{\bm{\rho}}$ onto the set of physical states $\mathbb{S}_{+}:=\{\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\bm{\rho}\succeq\bm{0},\operatorname{trace}(\bm{\rho})=1\}$. Denote by $P_{\mathbb{S}_{+}}$ the projection onto the set $\mathbb{S}_{+}$, which can be efficiently computed by projecting the eigenvalues onto a simplex [75]. Since the set $\mathbb{S}_{+}$ is convex, the corresponding projector is non-expansive, and hence
$\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-\bm{\rho}^{\star}\|_{F}$ $=$ $\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-P_{\mathbb{S}_{+}}(\bm{\rho}^{\star})\|_{F}$ (19)
$\leq$ $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq\epsilon,$
which implies that the projection step ensures that the state becomes physically valid while preserving or even improving the recovery guarantee.
Trace distance
The recovery guarantees discussed in this paper are established in the Frobenius norm (or Hilbert-Schmidt distance), which, while not a standard metric in quantum information, is widely adopted for analyzing MPO states and tensor networks [30, 76, 77]. This measure is particularly useful for understanding how errors scale with factors such as the sample size and the number of state copies in large-scale MPO states. Notably, although the size of the density matrix grows exponentially with $n$, the computational complexity of the Frobenius norm for an MPO state remains linear with respect to $n$ in [27]. In contrast, there is no known efficient method to compute the trace distance or fidelity between two MPO states. Therefore, the choice of using Frobenius norm to characterize the recovery errors for MPO states is natural from a computational perspective.
On the other hand, an issue with the Frobenius norm is that the state itself can potentially have an exponentially small Frobenius norm, particularly for high-rank states. However, this is not typically the case for many physical systems, such as those at low temperatures or near pure states. For example, the thermal/Gibbs state, $\bm{\rho}^{\star}=\frac{e^{-\bm{H}/T}}{\operatorname{trace}(e^{-\bm{H}/T})}$, where $\bm{H}$ denotes the Hamiltonian and $T$ the temperature (ignoring the Boltzmann constant) can be well approximated by a low-rank density matrix if $T$ is small. In addition, when $\bm{H}$ describes a one-dimensional spin model with short-range interactions, the resulting thermal state can often be approximated by an MPO. Furthermore, even for pure states with MPS representations (a special case of MPO states), there exist no rigorous sample complexity bounds that are efficient (i.e. the number of state copies is proportional to the number of independent parameters encoding the state).
To present a formal theorem formulated in terms of the trace distance for practical quantum states that are appropriately low-rank, we utilize the result in [78] to establish the connection between trace distance and the one based on Frobenius distance. In particular, suppose $\bm{\rho}^{\star}$ is approximately low-rank and can be expressed as $\bm{\rho}^{\star}=\bm{\rho}^{\star}_{s}+\bm{N}$, where $\bm{\rho}^{\star}_{s}$ is a rank-$s$ matrix [^7] Here $s$ denotes the matrix rank, while $\bm{r}=(r_{1},\dots,r_{n-1})$ represents MPO ranks. and $\bm{N}$ represents the small residual. Note that $\bm{N}=\bm{0}$ when the state is strictly low-rank.
Lemma 4.
Suppose $\bm{\rho}^{\star}=\bm{\rho}^{\star}_{s}+\bm{N}$, where $\bm{\rho}^{\star}_{s}$ is a rank-$s$ matrix. Then for any $\widehat{\bm{\rho}}\in\mathbb{C}^{d^{n}\times d^{n}}$, we have
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}\leq 2\sqrt{s}\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1}.$
Proof.
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}$ $\leq\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}_{s}\|_{1}+\|\bm{N}\|_{1}$
$\leq 2\sqrt{s}\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}_{s}\|_{F}+\|\bm{N}\|_{1}$
$=2\sqrt{s}(\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}+\bm{N}\|_{F})+\|\bm{N}\|_{1}$
$\leq 2\sqrt{s}\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1},$
where the second inequality follows from [78, Theorem 1]. ∎
Lemma 4 supports the utilization of Hilbert-Schmidt distance for practical quantum states, demonstrating that the trace distance can be upper bounded by the Hilbert-Schmidt distance. This property (with $\bm{N}=\bm{0}$) is utilized in [19] for obtaining guarantees in trace distance for low-rank states. Now, based on Lemma 4, we can extend Theorems 1 and 2 to trace distance. To unify results for both $2$-designs and $3$-designs, we define $\gamma_{t}(\bm{\rho}^{\star})=\begin{cases}\gamma(\bm{\rho}^{\star}),&t=2,\\ 1,&t>2.\end{cases}$
Corollary 1.
Suppose $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate $t$-design POVMs. Consider an MPO $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ that has MPO ranks $(r_{1},\dots,r_{n-1})$ and is approximately low-rank, i.e, $\bm{\rho}^{\star}=\bm{\rho}^{\star}_{s}+\bm{N}$. If we measure the state $M$ times using the POVM, then with high probability, the solution $\widehat{\bm{\rho}}$ of the constrained least squares (10) satisfies
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}$ $\lesssim\frac{d\overline{r}\sqrt{(1+\delta)ns\gamma_{t}(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}}$ (20)
$+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1}.$
As shown in Corollary 1, when the ground truth MPO state is approximately low-rank, the recovery error in terms of trace distance also exhibits a polynomial growth with respect to $n$.
Minimax lower bound
To further illustrate the sample efficiency of our result, we examine the minimax lower bound, a standard tool for assessing the tightness of the recovery upper bound in regression models. Note that the statistical noise $\bm{\eta}=\widehat{\bm{p}}-\bm{p}$ has mean zero and covariance matrix $\bm{\Sigma}$, where the elements are defined as ${\bm{\Sigma}}[i,j]=\begin{cases}\frac{p_{i}(1-p_{i})}{M},&i=j\\ -\frac{p_{i}p_{j}}{M},&i\neq j\end{cases}$. The covariance matrix $\bm{\Sigma}$ is approximately diagonal since the off-diagonal elements are expected to be significantly smaller than the diagonal elements; this is because $p_{i}$ is often very small due to the normalization property $\sum_{i}p_{i}=1$. In addition, when $M$ is sufficiently large, statistical noise can often be well-approximated by Gaussian noise. Consequently, Gaussian noise has been directly used to simulate statistical noise in the quantum state tomography [30]. Thus, to simplify the analysis, we temporarily assume the statistical noise $\bm{\eta}$ follows [^8] Here we assume covariance $\frac{1}{MK}{\bf I}$ to match the noise level: assume generic states where $\max_{i}{p_{i}}\leq 1/2$, then the expected squared norm of $\bm{\eta}$ is given by $\frac{1}{2M}\leq\operatorname{\mathbb{E}}\|\bm{\eta}\|_{2}^{2}=\operatorname{\mathbb{E}}\bigg{[}\sum_{k=1}^{K}\eta_{k}^{2}\bigg{]}=\sum_{k=1}^{K}\frac{p_{k}(1-p_{k})}{M}\leq\frac{1}{M},$ indicating that $\operatorname{\mathbb{E}}{\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}}$ has order of $1/M$. $\mathcal{N}({\bf 0},\frac{1}{MK}{\bf I})$, which satisfies $\operatorname{\mathbb{E}}\|\bm{\eta}\|_{2}^{2}=\frac{1}{M}$. Suppose $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate spherical $3$-design POVMs. Under these assumptions, it shares a similar setting to [79, Corollary 11]. By following a similar line of proof, we obtain the following minimax lower bound for any estimator $\widehat{\bm{\rho}}$
$\inf_{\widehat{\bm{\rho}}}\sup_{\bm{\rho}^{\star}\in\mathbb{X}_{\bm{r}}}\operatorname{\mathbb{E}}{\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}}\geq\Omega\bigg{(}\sqrt{\frac{nd^{2}\overline{r}^{2}}{M}}\bigg{)}.$ (21)
Compared to (21), the recovery upper bound in Theorem 2 is tight and optimal, up to logarithmic terms.
We now briefly discuss technical challenges in extending the above minimax lower bound to the statistical noise following multinomial distribution. Following [80, 19, 81], the first step is to construct a set of MPDO states $\{\bm{\rho}^{i}\}_{i=1}^{N}$ such that $\min_{i\neq j}\|\bm{\rho}^{i}-\bm{\rho}^{j}\|_{F}\geq\epsilon$. Then, we reformulate the recovery problem as a multiple hypothesis testing problem and apply the Fano’s Lemma to establish a lower bound on the minimax risk. This process involves computing an upper bound for the Kullback-Leibler (KL) divergence between two multinomial distributions of empirical measurements for $\bm{\rho}^{i}$ and $\bm{\rho}^{j}$, denoted by $\mathbb{Q}^{1}$ and $\mathbb{Q}^{2}$. Denoting the population measurements for $\bm{\rho}^{i}$ and $\bm{\rho}^{j}$ by $\{p_{k}\}$ and $\{q_{k}\}$, respectively, then the KL divergence [^9] Using the formulation for multinomial distributions, we have $\text{KL}(\mathbb{Q}^{1}||\mathbb{Q}^{2})=\sum_{k=1}^{K}\log\frac{p_{k}}{q_{k}}\sum_{f_{1},\dots,f_{K}}f_{k}\frac{M!}{\Pi_{k=1}^{K}f_{k}!}\\ \cdot\Pi_{k=1}^{K}q_{k}^{f_{k}}=\sum_{k=1}^{K}Mp_{k}\log\frac{p_{k}}{q_{k}}$ where the last equation follows $\sum_{f_{1},\dots,f_{K}}f_{k}\frac{M!}{\Pi_{k=1}^{K}f_{k}!}\Pi_{k=1}^{K}q_{k}^{f_{k}}=\operatorname{\mathbb{E}}{f_{k}}=Mp_{k}$. is given by $\text{KL}(\mathbb{Q}^{1}||\mathbb{Q}^{2})=\sum_{k=1}^{K}Mp_{k}\log\frac{p_{k}}{q_{k}}$. Bounding the KL divergence from above requires the probabilities $\{q_{k}\}$ to not be too small. For general states [81] or low-rank states [80], we can apply a unitary rotation on the constructed states $\{\bm{\rho}^{i}\}$, which remain general or low-rank states, to ensure that the probabilities $\{p_{i}\}$ and $\{q_{i}\}$ are approximately uniform. However, directly applying a global unitary rotation to MPDO states does not preserve the bond dimension of the original state, whereas applying a random local unitary rotation preserves the bond dimension but lacks effective statistical tools to guarantee approximately uniform probabilities. Despite this technical challenge, we conjecture that (21) remains valid for multinomial noise, similar to the necessary conditions for recovering general states [81] or low-rank states [80], which are proportional to the degrees of freedom of the target state. We leave a thorough study as future research. [^10] Recent work in [82] has established a necessary condition $\Omega(n^{\frac{1}{2}})$ for testing whether an unknown state is an MPS. However, it remains unclear whether this bound is tight, as the sufficient condition established in [82] still requires $O(nd\overline{r}^{2})$. Moreover, recovering an MPO state is expected to require much more state copies than testing whether an unknown state is an MPS.
4 Projected Gradient Descent with Guaranteed Convergence
In the literature of tensor recovery for tensor train (TT) format tensors, several iterative algorithms [83, 84, 85, 79, 86] have been proposed and analyzed from random measurements. While these algorithms can be applied to MPO by rewriting MPO as TT  [^11] To establish the equivalence, we first reshape $\bm{\rho}$ into an $n$-th order tensor $\mathcal{X}$ of size $d^{2}\times d^{2}\times\cdots\times d^{2}$, where each pair $(i_{\ell},j_{\ell})$ is mapped to a single index $s_{\ell}=i_{\ell}+d(j_{\ell}-1)$ for $\ell=1,\ldots,n$. Consequently, the $(s_{1},\ldots,s_{n})$-th element of $\mathcal{X}$ becomes $\mathcal{X}(s_{1},\dots,s_{n})=\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{s_{1}}\bm{X}_{2}^{s_{2}}\cdots\bm{X}_{n}^{s_{n}},$ where with abuse of notation we denote by $\bm{X}_{\ell}^{s_{\ell}}=\bm{X}_{\ell}^{i_{\ell},j_{\ell}}$., the analysis has predominantly focused on real-valued tensors with either clear measurements or corrupted by Gaussian noise. Thus, the guarantees cannot be directly applied to QST involving complex-valued MPOs and statistical noise induced by the empirical measurements that follow a multinomial distribution. In this section, we propose an iterative optimization algorithm and study its convergence for solving the constrained least-squares problem (10) that minimizes the discrepancy between the empirical measurements $\widehat{\bm{p}}$ and the linear mapping of the estimated MPO $\bm{\rho}$.
We begin by reiterating the loss function in (10):
$\min_{\bm{\rho}\in\mathbb{X}_{\bm{r}}}g(\bm{\rho})=\|\mathcal{A}(\bm{\rho})-\widehat{\bm{p}}\|_{2}^{2}.$ (22)
We solve this constrained optimization problem by a projected gradient descent (PGD) that iteratively updates
$\bm{\rho}^{(\tau+1)}=\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})),$ (23)
where $\mu$ is the step size, $\nabla g(\bm{\rho}^{(\tau)})=\sum_{k=1}^{K}(\langle\bm{A}_{k},\bm{\rho}^{(\tau)}\rangle-\widehat{\bm{p}})\bm{A}_{k}$ is the Wirtinger gradient  [^12] Note that $g(\bm{\rho})$ is non-holomorphic (i.e., not complex differentiable), which poses challenges for the development and analysis of standard gradient-based optimization algorithms. Fortunately, we can adopt Wirtinger gradient to extend the classical gradient concept to functions of complex variables [87]. In our context, we first parameterize the objective function as $g(\bm{\rho},\bm{\rho}^{*})=\sum_{k=1}^{K}(\langle\bm{A}_{k}^{*},\bm{\rho}^{*}\rangle-\widehat{p}_{K}^{*})(\langle\bm{A}_{k},\bm{\rho}\rangle-\widehat{p}_{K})$, and then compute the Wirtinger gradient $\nabla_{\bm{\rho}^{*}}g(\bm{\rho},{\bm{\rho}^{*}})$ by treating $\bm{\rho}$ and $\bm{\rho}^{*}$ as two independent variables. We can then use $\nabla_{\bm{\rho}^{*}}g(\bm{\rho},{\bm{\rho}^{*}})$ as the gradient in standard gradient-based optimization algorithms. To simplify the notation, we denote the Wirtinger gradient $\nabla_{\bm{\rho}^{*}}g(\bm{\rho},{\bm{\rho}^{*}})$ simply as $\nabla g(\bm{\rho})$., and $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)$ is the projection onto the set $\mathbb{X}_{\bm{r}}$.
Since computing the optimal projection onto the set of TT format tensors is already known as NP-hard [41], computing the optimal projection onto the set $\mathbb{X}_{\bm{r}}$ is challenging even ignoring the constraint on the trace. Here, we use a two-step approximate projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)=\mathcal{P}_{\operatorname{trace}=1}(\text{SVD}_{\bm{r}}^{tt}(\cdot))$, where $\text{SVD}_{\bm{r}}^{tt}(\cdot)$ denotes the TT-SVD operation [27] that can compute an MPO (may not have trace 1) approximation for a given density matrix  [^13] The output $\text{SVD}_{\bm{r}}^{tt}(\bm{A})$ is always Hermitian if $\bm{A}$ is Hermitian., and $\mathcal{P}_{\operatorname{trace}=1}(\cdot)$ represents the projection onto the convex set $\{\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\operatorname{trace}(\bm{\rho})=1\}$ to have unit trace  [^14] This can be achieved by computing the eigendecomposition and subsequently projecting the eigenvalues onto the simplex [75]. While this process may increase the MPO ranks, our observations indicate that this increase is often minimal or does not occur. However, this process (particularly computing the eigendecomposition) could be computationally expensive for large quantum states. Alternatively, in our implementation, we simply use $\mathcal{P}_{\operatorname{trace}=1}(\bm{\rho})=\bm{\rho}/\operatorname{trace}(\bm{\rho})$ which is computationally efficient while simultaneously preserving the MPO structure. Thus, in the analysis, we will assume the projection $\mathcal{P}_{\operatorname{trace}=1}(\cdot)$ also preserves the MPO structure of the input.. While the TT-SVD operation $\text{SVD}_{\bm{r}}^{tt}(\cdot)$ may not produce an optimal approximation, it can be approved to be sub-optimal. In particular, for any density matrix $\bm{\rho}$, we have [27] $\|\text{SVD}_{\bm{r}}^{tt}(\bm{\rho})-\bm{\rho}||_{F}^{2}\leq(n-1)\min_{\bm{\rho}^{\prime}\in\mathbb{X}_{\bm{r}}}\|\bm{\rho}^{\prime}-\bm{\rho}\|_{F}^{2}$. Moreover, the following result provides an improved bound when $\bm{\rho}$ can be approximated by an MPO.
Lemma 5.
([88, Lemma 26]) Let $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ be an MPO with MPO ranks $(r_{1},\dots,r_{n-1})$. Denote by $\underline{\sigma}({\bm{\rho}^{\star}})$ the smallest TT singular value of $\bm{\rho}^{\star}$ [^15] Specifically, we reshape $\bm{\rho}$ into an $n$-th order tensor $\mathcal{X}$ of size $d^{2}\times d^{2}\times\cdots\times d^{2}$ and define the $\ell$-th unfolding matrix of tensor $\mathcal{X}$ as $\bm{X}^{\langle\ell\rangle}\in\mathbb{C}^{d^{2\ell}\times d^{2n-2\ell}}$, where the $(s_{1}\cdots s_{\ell},s_{\ell+1}\cdots s_{n})$-th element of $\bm{X}^{\langle\ell\rangle}$ is given by $\bm{X}^{\langle\ell\rangle}(s_{1}\cdots s_{\ell},s_{\ell+1}\cdots s_{n})=\mathcal{X}(s_{1},\dots,s_{n})$. With the $\ell$-th unfolding matrix $\mathcal{X}^{\langle\ell\rangle}$ and the MPO ranks, we can obtain its smallest TT singular value $\underline{\sigma}(\mathcal{X})=\min_{\ell=1}^{n-1}\sigma_{r_{\ell}}(\mathcal{X}^{\langle\ell\rangle})$.. Then for any perturbation $\bm{E}\in\mathbb{C}^{d^{n}\times d^{n}}$ with $C_{n}||\bm{E}||_{F}\leq\underline{\sigma}({\bm{\rho}^{\star}})$ for some constant $C_{n}\geq 500n$, we have
$\|\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{\star}+\bm{E})-\bm{\rho}^{\star}\|_{F}^{2}\leq\|\bm{E}||_{F}^{2}+\frac{600n||\bm{E}\|_{F}^{3}}{\underline{\sigma}({\bm{\rho}^{\star}})}.$
Local convergence of PGD
We now provide convergence analysis for the PGD in (23). Without loss of generality, our analysis will be presented for the general $\delta$-approximate $t$-design POVM. To unify results for both $2$-designs and $3$-designs, recall the definition $\gamma_{t}(\bm{\rho}^{\star})=\begin{cases}\gamma(\bm{\rho}^{\star}),&t=2,\\ 1,&t>2.\end{cases}$
Theorem 3 (Local linear convergence of PGD).
Suppose that $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate $t$-design POVMs, which is used to measure an MPO state $\bm{\rho}^{\star}$ $M$ times. Given an initialization $\bm{\rho}^{(0)}$ satisfying
$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}<\frac{\underline{\sigma}({\bm{\rho}^{\star}})(d-1)(1-\delta)^{2}}{600n(1+\delta^{2}+(4d-2)\delta)},$ (24)
then with probability at least $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, the PGD in (23) with step size $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ converges at a linear rate and
$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg{(}\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$ (25)
for sufficiently large $\tau$.
The proof is given in Appendix E. Theorem 4 provides a linear convergence analysis of the PGD algorithm along with its error characterization. Moreover, the recovery error presented in (27) aligns closely with the results in Theorems 1 and 2. Since the problem (22) is nonconvex (as the MPO set is nonconvex), Theorem 4 requires an initialization close to the target state that satisfies (24). [^16] The right-hand side of (24) depends on $\underline{\sigma}({\bm{\rho}^{\star}})$, which could potentially be exponentially small for high-rank states. However, this is not the case for many practical states, such as those at low temperatures or systems close to pure states. However, as will be illustrated in Section 5, our experiments show that even a randomly initialized PGD can efficiently converge to a stable recovery error. To relax the requirement on the initialization, the following result establishes another convergence guarantee with an additional requirement on the projection.
Theorem 4 (Linear convergence of PGD).
Suppose that $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate $t$-design POVMs, which is used to measure an MPO state $\bm{\rho}^{\star}$ $M$ times. Given an initialization $\bm{\rho}^{(0)}$, suppose the PGD in (23) uses step size $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ and that the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}$ satisfies
$\|\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F}^{2}\leq(1+b)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2},$
$\text{with}\ b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}.$ (26)
Then with probability at least $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, the PGD converges at a linear rate and
$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg{(}\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$ (27)
for sufficiently large $\tau$.
The proof is given in Appendix E. Although the two-step approximate projection may lack rigorous proof of satisfying (4), we will demonstrate the convergence of the PGD algorithm with a random initialization, implying that $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)=\mathcal{P}_{\operatorname{trace}=1}(\text{SVD}_{\bm{r}}^{tt}(\cdot))$ may satisfy the assumption in (4).
Spectral initialization
We now discuss some potential initialization strategies for the PGD. A common approach to generating an appropriate initialization for signal estimation in nonconvex scenarios in the spectral initialization approach. It has been widely employed for various inverse problems [89], such as phase retrieval [90, 91], low-rank matrix recovery [92, 93], and structured tensor recovery [94, 95, 79]. In our context, this entails simply computing an MPO approximation to $\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\widehat{p}_{k}\bm{A}_{k}$, where $\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\widehat{p}_{k}\bm{A}_{k}$ is equivalent to apply the adjoint operator of $\mathcal{A}$ on the empirical measurements $\widehat{\bm{p}}$ with a scaling factor $K(d^{n}+1)/d^{n}$ that is used to balance the energy according to (2). Specifically, by using the same projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}$ in the PGD, we can compute an initialization via
$\bm{\rho}^{(0)}=\mathcal{P}_{\mathbb{X}_{\bm{r}}}\bigg{(}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\widehat{p}_{k}\bm{A}_{k}\bigg{)}.$ (28)
The following result ensures that the initialization $\bm{\rho}^{(0)}$ is close to $\bm{\rho}^{\star}$.
Theorem 5.
Suppose that $\{\bm{A}_{1},\ldots,\bm{A}_{K}\}$ form a set of $\delta$-approximate $t$-design POVMs, which is used to measure an MPO state $\bm{\rho}^{\star}$ $M$ times. By utilizing the spectral initialization in (28), with probability $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, we have
[…]
$\cdot O\bigg{(}\frac{d\overline{r}\sqrt{(1+\delta)n\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}+\delta\|\bm{\rho}^{\star}\|_{F}\bigg{)}.$
Its proof is given in Appendix F. The scaling term $\sqrt{n-1}$ arises due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)=\mathcal{P}_{\operatorname{trace}=1}(\text{SVD}_{\bm{r}}^{tt}(\cdot))$. Recall that $\gamma_{t}(\bm{\rho}^{\star})=1$ for all $\bm{\rho}^{\star}$ when spherical 3-design POVMs (i.e., $t\geq 3$), and $\gamma_{t}(\bm{\rho}^{\star})=\gamma(\bm{\rho}^{\star})$ for spherical 2-design POVMs (i.e., $t=2$). Theorem 5 shows that when $\gamma_{t}(\bm{\rho}^{\star})$ is small—true for all MPOs when using $3$-designs and for those having relatively uniformed probability distribution with 2-design POVMs—the spectral initialization provides a relatively good estimation with a polynomially large $M$. In particular, considering exact $t$-design POVMs or cases where $\delta$ is small, (29) ensures that $\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}$ scales in ther order of $\frac{n^{2}d^{2}\overline{r}^{2}}{M}$ when ignoring the log terms and $\gamma_{t}(\bm{\rho}^{\star})$. Note that the requirements on the number of measurements are slightly more stringent than those in Theorem 1 and Theorem 2 due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)$. However, as will be illustrated in Section 5, our numerical experiments show that even a randomly initialized PGD can very efficiently find an estimate of the target state. A theoretical explanation for this phenomenon will be the focus of future research.
Projected stochastic gradient descent
The primary computational complexity of the PGD (23) arises from computing the gradient, which involves an exponentially large number ($K$) of inner product operations between $\bm{A}_{k}$ and $\bm{\rho}^{(\tau)}$, with a complexity of $O(Kd^{2n})$ by naive computation. When the POVMs $\{\bm{A}_{k}\}_{k=1}^{K}$ also have MPO representations, such as the local measurements as detailed in the subsequent section, the complexity of each inner product can be reduced to $O(n{\overline{r}}^{3})$ from $O(d^{2n})$ by using the dot product method [27] for two MPOs. Similarly, the gradient update $\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})$ in (23) can also be represented in the MPO form rather than as a full density matrix, allowing us to utilize TT-rounding to efficiently compute TT-SVD with polynomial computational complexity in $n$ [27].
The issue of computing all the $K$ inner products in one iteration can be addressed by using stochastic methods [96]. In our context, exploiting the fact that the empirical probabilities $\{\widehat{p}_{i}\}$ are very sparse (at most $M$ of them are nonzero) when the measured time $M$ is much smaller than $K$, we can utilize a customized stochastic method. Specifically, unlike standard approaches that either randomly or sequentially select a subset of samples for computing gradient in each iteration, we can use the nonzero measurements more frequently. For instance, each time (often called epoch in machine learning), we first select a subset (say $N$) of empirical probabilities that include all the nonzero ones and a few numbers of zero ones (which can be either randomly selected or selected in a sequential), and then perform $N/B$ iterations of PSGD with each iteration using $B$ measurements selected sequentially form this subset.
5 Numerical Experiments with Local IC-POVMs
To our best knowledge, there are no known efficient implementations of SIC-POVM or spherical $t$-designs with $t\geq 2$ for a many-qubit system using local quantum circuits available in current or near-future quantum hardware. It is not surprising that the sample-efficient measurement setting for recovering an MPO state is not efficiently implementable, as this is also the case for the QST of generic quantum states [19, 16, 17]. To make our results more experimentally relevant, in this section we consider a different type of IC-POVM that is efficiently implementable. Specifically, we consider the following local SIC-POVM $\{\bm{B}_{i}\}_{i=1}^{4}$ for each qubit
$\bigg{\{}\begin{bmatrix}\frac{1}{2}&0\\ 0&0\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}\\ \frac{\sqrt{2}}{6}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{2\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{2\pi}{3}}&\frac{1}{6}\end{bmatrix},$
$\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{4\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{4\pi}{3}}&\frac{1}{6}\end{bmatrix}\bigg{\}}.$ (30)
We can then generate the set of IC-POVM for an $n$-qubit system as $\{\bm{A}_{i}\}_{i=1}^{4^{n}}=\{\bm{B}_{i_{1}}\otimes\cdots\otimes\bm{B}_{i_{n}}\}_{i_{1},\dots,i_{n}}$. Importantly, such IC-POVM has been recently demonstrated experimentally using trapped ions in a scalable way [97] and can be implemented in most quantum computing hardware platforms using either ancilla qubits or extra low-energy states in the physical carriers of the qubits.
Figure 1: Illustration of (a) recovery error for different initialization methods, $\overline{r}$ and $n$ with $M=3000$; (b) convergence rate of PGD for different $n$, spectral initialization (dotted lines), and random initialization (solid lines) with $M=3000$ and $\overline{r}=4$; (c) recovery error with $n=5$ for different initialization methods, $\overline{r}$ and $M$; (d) $p_{\text{max}}$ for the MPDO.
While our main results (Theorems 1 and 2) do not apply to such a measurement setting, we show that certain MPO states can be recovered using this IC-POVM using a number of state copies only linear in $n$ and the PGD and PSGD algorithms developed in Section 3. A more systematic study on what types of MPO states can be recovered efficiently using the local IC-POVM can be found in [98], using a different optimization algorithm based on maximal likelihood estimation (MLE). As an example, here we consider a random matrix product density operator (MPDO) state [68] of the following form (1), i.e, the $(i_{1}\cdots i_{n},j_{1}\cdots j_{n})$-th entry of $\bm{\rho}^{\star}$ can be generated by
$\bm{\rho}^{\star}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})={\bm{X}_{1}^{\star}}^{i_{1},j_{1}}{\bm{X}_{2}^{\star}}^{i_{2},j_{2}}\cdots{\bm{X}_{n}^{\star}}^{i_{n},j_{n}}.$
To ensure $\bm{\rho}^{\star}$ PSD, we generate each matrix ${\bm{X}_{\ell}^{\star}}^{i_{\ell},j_{\ell}}$ as [68]
${\bm{X}_{\ell}^{\star}}^{i_{\ell},j_{\ell}}=\sum_{a_{\ell}=1}^{K_{\ell}}{\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}\otimes({\bm{A}_{\ell}^{\star}}^{j_{\ell},a_{\ell}})^{*},$ (31)
where the symbol * represents the complex conjugate, $\otimes$ denotes the Kronecker product, $K_{\ell}$ is an arbitrary positive integer that controls purity of the generated state, and each matrix ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ has dimension $\kappa\times\kappa$, except for ${\bm{A}_{1}^{\star}}^{i_{1},a_{1}}$ and ${\bm{A}_{n}^{\star}}^{i_{n},a_{n}}$, which are of dimension $1\times\kappa$ and $\kappa\times 1$ respectively. Each entry of ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ is i.i.d. randomly generated with both its real and imaginary parts drawn uniformly from the interval $[-1,1]$. In the experiments, we set $K_{\ell}=10$ to ensure the state is sufficiently mixed; using $K_{\ell}=1$ would generate a pure state. The generated $\bm{\rho}^{\star}$ is an MPO with a bond dimension of $\overline{r}=\kappa^{2}$ and is always PSD. To be a physical state, $\bm{\rho}^{\star}$ also needs to have unit trace, which can be done by calculating
$\operatorname{trace}(\bm{\rho}^{\star})=({\bm{X}_{1}^{\star}}^{1,1}+{\bm{X}_{1}^{\star}}^{2,2})\cdots({\bm{X}_{n}^{\star}}^{1,1}+{\bm{X}_{n}^{\star}}^{2,2}),$
and then dividing each matrix ${\bm{X}_{\ell}^{\star}}^{i_{\ell},j_{\ell}}$ by $(\operatorname{trace}(\bm{\rho}^{\star}))^{\frac{1}{n}}$.
In order to validate the theoretical analysis presented in the preceding sections, we apply PGD and PSGD with spectral initialization across various measurements. To further mitigate the high computational complexity of the spectral method, we will also use a random initialization generated by a random MPDO as the starting point for both PGD and PSGD. This approach is motivated by our observation that random initialization, when the target is a random MPDO, can effectively serve as a convergence-guaranteeing initialization.
Figure 2: (a)Convergence comparison of PSGD with random initialization for various $n$; (b) Performance comparison between PGD and PSGD with random and spectral initializations.
Performance of PGD
In the first set of experiments, we set $M=3000$ for empirical measurements. Since we will perform experiments for different number of qubits, we employ PGD with diminishing step sizes $\mu_{\tau}=\mu_{0}\cdot 2^{n}\cdot\lambda^{\tau}$ (with $\lambda=0.9$), which is found to provide stable performance across varying $n$. When a random initialization is used, we set $\mu_{0}=\frac{5}{4}$ for $\overline{r}=1$ and $\mu_{0}=\frac{5}{8}$ for $\overline{r}=4$. As for spectral initialization, $\mu_{0}$ is set to $\frac{5}{8}$ for $\overline{r}=1$ and $\frac{5}{16}$ for $\overline{r}=4$. Figure 1 depicts the recovery error as a function of $n$ for various $\overline{r}$ and initialization methods. As anticipated, the recovery error tends to rise with an increase in $\overline{r}$, yet the PGD algorithm displays stability across all cases. Notably, we observe that the recovery error increases only polynomially rather than exponentially with respect to $n$, aligning with the findings in Theorems 1 and 2. Furthermore, Figure 1 illustrates the convergence behavior of PGD with both spectral and random initializations. As $n$ increases, the quality of initial values from spectral initialization deteriorates, aligning with the theoretical analysis presented in Theorem 5. Conversely, although the recovery error with random initialization is slightly worse than that of the spectral method, it offers a reliable starting point for stable recovery while avoiding the high computational complexity associated with the spectral method. Subsequently, we evaluate the recovery performance across varying values of $M$, selecting $\mu_{0}=\frac{5}{32}$ for random initialization and $\mu_{0}=\frac{1}{16}$ for spectral initialization, with $\lambda=1$. In Figure 1, we observe that the recovery performance of PGD with spectral and random initializations is comparable, with improved recovery error as $M$ increases. Finally, Figure 1 reveals that the numerator of the average maximum probability is polynomial, ensuring the effectiveness of the assumption.
Performance comparison between PGD and PSGD
In the second experiment, we compare the performance of PGD and PSGD in estimating the MPDO. We set $M=3000$ and $\overline{r}=4$. We use the same setting for PGD in Figure 1 and use diminishing step sizes $\mu_{\tau}=\frac{5}{4}\times 2^{n}\times 0.9^{\tau}$ for random initialization and $\mu_{\tau}=10\times 0.9^{\tau}$ for spectral initialization. Additionally, we choose $N=10\times 4n\overline{r}^{2}=640n$ measurements in one epoch and $B=32$ in each iteration of one epoch. Figure 2 provides an analysis of the convergence properties of PSGD with random initialization, demonstrating that this achieves a linear convergence rate and exhibits stable recovery errors as $n$ increases. Furthermore, due to its polynomial computational complexity, PSGD with random initialization is viable for systems with up to 14 qubits. In Figure 2, we observe that PSGD with random initialization delivers performance comparable to that achieved with spectral initialization. Notably, although PSGD with polynomial measurements incurs a marginally higher recovery error compared to PGD, it nonetheless achieves a stable recovery error that grows polynomially with $n$.
To further elucidate the computational complexity of the two methods, we summarize the computational cost for a single iteration of PGD and one epoch of PSGD under local measurements as follows:
[…]
6 Conclusion
This paper focuses on the sample complexity bounds for recovering structured quantum states in one-dimension that are represented as matrix product operators (MPOs) with finite bond dimensions. The study begins by establishing theoretical bounds on the accuracy of a constrained least-squares estimator for MPO state recovery, utilizing empirical measurements from a class of Informationally Complete Positive Operator-Valued Measure (IC-POVM). The results indicate that assuming probabilities measured by one POVM are approximately uniform for SIC-POVM or $2$-designs, or without any assumption for $t$-designs ($t\geq 3$), a stable recovery guarantee requires the number of state copies to be only linear in the the number of qudits $n$ and is thus sample efficient. Additionally, the paper introduces a provable projected gradient descent (PGD) algorithm. It demonstrates that, with appropriate initialization achievable through spectral initialization, PGD converges to the neighborhood of the target state at a linear rate and achieves an error bound scaling only linear in $n$. These results address fundamental questions on whether quantum states with efficient classical representations can be recovered efficiently. While the measurement settings used in our sample complexity bounds are not efficiently implementable, we show that certain MPO states may be recovered efficiently using a local IC-POVM that can be experimentally achieved on current quantum hardware. It remains an open question what class of MPO states, or other structured quantum states, can be recovered not only using an efficient number of state copies, but also using an efficiently implementable measurement setting [98].
Acknowledgment
We acknowledge funding support from NSF Grants No. CCF-1839232, PHY-2112893, CCF-2106834, CCF-2241298 and ECCS-2409701 as well as the W. M. Keck Foundation. We thank the Ohio Supercomputer Center for providing the computational resources needed in carrying out this work.
Appendix A Proof of Lemma 1
Proof.
According to [99], any Hermitian matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$ can be represented in terms of elements of the dual basis as
$\bm{\rho}=\sum_{k=1}^{d^{2n}}\langle\bm{A}_{k},\bm{\rho}\rangle\widetilde{\bm{A}}_{k}.$ (32)
Here, the dual basis $\{\widetilde{\bm{A}}_{k}\}$ is a basis such that $\langle\bm{A}_{k},\widetilde{\bm{A}}_{j}\rangle=\begin{cases}1,&k=j\\ 0,&k\neq j\end{cases}$. For a general SIC-POVM $\{\bm{A}_{k}\}$, the dual basis is comprised by operators [100]
$\widetilde{\bm{A}}_{k}=d^{n}(d^{n}+1)\bm{A}_{k}-{\bf I}_{d^{n}}.$ (33)
Using $\bm{\rho}$ in (32), we have
$\operatorname{trace}(\bm{\rho}^{2})$ $=$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\sum_{j\neq k}\langle\bm{A}_{j},\bm{\rho}\rangle\langle\bm{A}_{k},\bm{\rho}\rangle$ (34)
$=$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\bigg{(}\bigg{(}\sum_{i=1}^{d^{2n}}\langle\bm{A}_{i},\bm{\rho}\rangle\bigg{)}^{2}-\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\bigg{)}$
$=$ $d^{n}(d^{n}+1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-(\operatorname{trace}(\bm{\rho}))^{2},$
where the first line follows (22) in [99, Proposition 1]. In addition, due to $\bm{\rho}=\bm{\rho}^{\dagger}$, we have $\operatorname{trace}(\bm{\rho}^{2})=\operatorname{trace}(\bm{\rho}^{\dagger}\bm{\rho})=\|\bm{\rho}\|_{F}^{2}$. This completes the proof. ∎
Appendix B Proof of LABEL:{l2_norm_of_approximate_2-designs_RIP}
Proof.
According to [40, Theorem 1] and Lemma 6, we have
[…]
where $\bm{S}$ is the swap operator. Taking the inner product between (35) and $\frac{d^{2n}}{K}\bm{\rho}^{\otimes 2}$ we get
[…]
$=$ $\frac{d^{n}((\operatorname{trace}(\bm{\rho}))^{2}+\operatorname{trace}(\bm{\rho}^{2}))}{K(d^{n}+1)}$
$=$ $\frac{d^{n}((\operatorname{trace}(\bm{\rho}))^{2}+\|\bm{\rho}\|_{F}^{2})}{K(d^{n}+1)},$
where the penultimate line follows $\operatorname{trace}(\bm{S}\bm{\rho}^{\otimes 2})=\operatorname{trace}(\bm{\rho}^{2})$ [15, Lemma 17] for any Hermitian matrix $\bm{\rho}$. Based on $\operatorname{trace}((\bm{w}_{k}\bm{w}_{k}^{\dagger})^{\otimes 2}(\bm{\rho}^{\otimes 2}))=\operatorname{trace}((\bm{w}_{k}\bm{w}_{k}^{\dagger}\bm{\rho})\otimes(\bm{w}_{k}\bm{w}_{k}^{\dagger}\bm{\rho}))=(\operatorname{trace}(\bm{w}_{k}\bm{w}_{k}^{\dagger}\bm{\rho}))^{2}$, this completes the proof. ∎
Appendix C Proof of Theorem 1
Proof.
Now, we start by upper-bounding $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$. Towards that goal, we first rewrite this term as
$\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle=\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle\leq\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle,$ (37)
where we denote by $\overline{\mathbb{X}}_{\bm{r}}$ the normalized set of MPOs with $\bm{r}=[r_{1},\dots,r_{n-1}]$:
$\begin{split}\overline{\mathbb{X}}_{\bm{r}}=\Big{\{}&\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\ \bm{\rho}=\bm{\rho}^{\dagger},\|\bm{\rho}\|_{F}\leq 1,\operatorname{trace}(\bm{\rho})=0,\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},\\ &\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}},\ell\in[n-1],r_{0}=r_{n}=1\Big{\}}.\end{split}$ (38)
It is important to highlight that in comparison to $\mathbb{X}_{\bm{r}}$ in (11), $\overline{\mathbb{X}}_{\bm{r}}$ additionally includes $\|\bm{\rho}\|_{F}\leq 1$ and $\operatorname{trace}(\bm{\rho})=0$.
The rest of the proof is to bound $\max_{\bm{\rho}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle$, which will be achieved by using a covering argument. First, when conditioned on $\bm{A}_{k}$, we consider any fixed value of $\bm{\rho}^{(p)}=[\bm{X}_{1}^{(p_{1})},\dots,\bm{X}_{n}^{(p_{n})}]\in\overline{\mathbb{X}}_{2\bm{r}}$ and according to [33, Appendix B-A], construct a $\xi$-net $\{\bm{\rho}^{(1)},\ldots,\bm{\rho}^{(N_{1}\cdots N_{n})}\}=\{[\bm{X}_{1}^{(1)},\dots,\bm{X}_{n}^{(1)}],\dots,[\bm{X}_{1}^{(N_{1})},\dots,\bm{X}_{n}^{(N_{n})}]\}$ such that $\sup_{\bm{X}_{\ell}:\|\bm{X}_{\ell}\|_{F}\leq 1}\leavevmode\nobreak\ \min_{p_{\ell}\leq N_{\ell}}\|\bm{X}_{\ell}-\bm{X}_{\ell}^{(p_{\ell})}\|_{F}\leq\xi,\ell=1,\dots,n$ with covering number $\Pi_{\ell=1}^{n}N_{\ell}\leq(\frac{4+\xi}{\xi})^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}$.
Then we apply Lemma 3 to establish a concentration inequality for the expression $\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle$ as follows:
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle\geq t\right)$$\leq$ $e^{-\frac{Mt}{4\max_{k}|\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle|}\min\bigg{\{}1,\frac{\max_{k}|\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle|t}{4\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}p_{k}}\bigg{\}}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}p_{k}}}$ (39)
$\leq$ $e^{-\frac{Mt^{2}}{16\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}p_{k}}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}p_{k}}},$
where without loss of generality, we assume that $\frac{\max_{k}|\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle|t}{4\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}p_{k}}\leq 1$ in the last line.
Since $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}\leq\frac{(1+\delta)d^{n}}{K(d^{n}+1)}$ from Lemma 2 and $\max_{k}p_{k}=\frac{\gamma(\bm{\rho}^{\star})}{K}$, we can derive $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle^{2}p_{k}\leq\frac{(1+\delta)d^{n}\gamma(\bm{\rho}^{\star})}{K^{2}(d^{n}+1)}$ and thus, have
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle\geq t\right)\leq 2e^{-\frac{K^{2}(d^{n}+1)Mt^{2}}{16(1+\delta)d^{n}\gamma(\bm{\rho}^{\star})}}.$ (40)
According to the same analysis of [33, eq.(110)] with the covering number $(\frac{4+\xi}{\xi})^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}$, we can derive
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle\geq t\right)$$\leq$ $\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle\geq\frac{t}{2}\right)$ (41)
$\leq$ $\bigg{(}\frac{4+\xi}{\xi}\bigg{)}^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}e^{-\frac{K^{2}(d^{n}+1)Mt^{2}}{16(1+\delta)d^{n}\gamma(\bm{\rho}^{\star})}+\log 2}$
$\leq$ $e^{-\frac{K^{2}(d^{n}+1)Mt^{2}}{16(1+\delta)d^{n}\gamma(\bm{\rho}^{\star})}+Cnd^{2}\overline{r}^{2}\log n+\log 2},$
where the first equation follows [33, eq.(89)] with $\xi=\frac{1}{2n}$, $\overline{r}=\max_{i}r_{i}$, and $C$ is a universal constant in the last line. By taking $t=\frac{c_{1}d\overline{r}\sqrt{(1+\delta)nd^{n}\gamma(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}$ in the above equation, we further obtain
$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle\leq\frac{c_{1}d\overline{r}\sqrt{(1+\delta)nd^{n}\gamma(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\right)\geq 1-e^{-c_{2}nd^{2}\overline{r}^{2}\log n},$ (42)
where $c_{1}$ and $c_{2}$ are constants.
Hence, with (37), we can obtain
$\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle\leq\frac{c_{1}d\overline{r}\sqrt{(1+\delta)nd^{n}\gamma(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}.$ (43)
Combining (13) and (14), with probability $1-e^{-c_{2}nd^{2}\overline{r}^{2}\log n}$, we arrive at
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq\frac{c_{3}d\overline{r}\sqrt{(1+\delta)n\gamma(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}},$ (44)
where $c_{3}$ is a constant. ∎
Appendix D Proof of Theorem 2
Proof.
Firstly, we derive the upper bound of $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ for $\bm{\rho},\bm{\rho}^{\star}\in\mathbb{X}_{\bm{r}}$ in (11). Leveraging the ties of $\delta$-approximate $3$-designs [67, 40] and Lemma 6, we can get:
$\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ $=$ $\sum_{k=1}^{K}\langle\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}\langle\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger},\bm{\rho}^{\star}\rangle$ (45)
$=$ $\sum_{k=1}^{K}\frac{8^{n}}{K^{3}}\operatorname{trace}((\bm{w}_{k}\bm{w}_{k}^{\dagger})^{\otimes 3}(\bm{\rho}-\bm{\rho}^{\star})\otimes(\bm{\rho}-\bm{\rho}^{\star})\otimes\bm{\rho}^{\star})$
$\leq$ $\frac{d^{3n}}{K^{2}}\frac{6(1+\delta)}{(d^{n}+2)(d^{n}+1)d^{n}}\operatorname{trace}(\mathcal{P}_{\text{sym}^{\otimes 3}}(\bm{\rho}-\bm{\rho}^{\star})\otimes(\bm{\rho}-\bm{\rho}^{\star})\otimes\bm{\rho}^{\star}),$
where $\mathcal{P}_{\text{sym}^{\otimes 3}}$, defined in [40, Lemma 3], represents the projector onto the symmetric subspace. Additionally, based on [63, Lemma 7] and [101, eq. (322)], we have
$\operatorname{trace}(\mathcal{P}_{\text{sym}^{\otimes 3}}(\bm{\rho}-\bm{\rho}^{\star})\otimes(\bm{\rho}-\bm{\rho}^{\star})\otimes\bm{\rho}^{\star})$ (46)
$=$ $\frac{1}{6}\bigg{(}(\operatorname{trace}(\bm{\rho}-\bm{\rho}^{\star}))^{2}\operatorname{trace}(\bm{\rho}^{\star})+\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})^{2})\operatorname{trace}(\bm{\rho}^{\star})$
$+2\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})\bm{\rho}^{\star})\operatorname{trace}(\bm{\rho}-\bm{\rho}^{\star})+2\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})^{2}\bm{\rho}^{\star})\bigg{)}$
$=$ $\frac{1}{6}\|\bm{\rho}-\bm{\rho}^{\star}\|_{F}^{2}+\frac{1}{3}\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})^{2}\bm{\rho}^{\star}).$
Note that this can also be directly obtained from [81, eq. (S36)] when $\operatorname{trace}(\bm{\rho}-\bm{\rho}^{\star})=0$. Consequently, we arrive at $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}\leq O((1+\delta)\frac{\|\bm{\rho}-\bm{\rho}^{\star}\|_{F}^{2}+\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})^{2}\bm{\rho}^{\star})}{K^{2}})$. Given that $\bm{\rho}-\bm{\rho}^{\star}$ is Hermitian, we can ensure $(\bm{\rho}-\bm{\rho}^{\star})^{2}=(\bm{\rho}-\bm{\rho}^{\star})(\bm{\rho}-\bm{\rho}^{\star})^{\dagger}$ is PSD. Building upon [102, Theorem 1] using two PSD matrices $(\bm{\rho}-\bm{\rho}^{\star})^{2}$ and $\bm{\rho}^{\star}$, we further deduce $\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})^{2}\bm{\rho}^{\star})\leq\operatorname{trace}((\bm{\rho}-\bm{\rho}^{\star})^{2})\operatorname{trace}(\bm{\rho}^{\star})=\|\bm{\rho}-\bm{\rho}^{\star}\|_{F}^{2}$. Ultimately, we obtain
$\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}\leq O\bigg{(}\frac{(1+\delta)\|\bm{\rho}-\bm{\rho}^{\star}\|_{F}^{2}}{K^{2}}\bigg{)}.$ (47)
Applying the same analysis as in Appendix C, we can derive
$\operatorname{\mathbb{P}}\left(\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle\leq\frac{c_{1}d\overline{r}\sqrt{(1+\delta)n\log n}}{K\sqrt{M}}\right)\geq 1-e^{-c_{2}nd^{2}\overline{r}^{2}\log n},$ (48)
where $\overline{\mathbb{X}}_{2\bm{r}}$ is defined in (38) and $c_{1}$, $c_{2}$ are constants.
Combining (13), (14) and (48), we can conclude:
$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq O\bigg{(}\frac{d\overline{r}\sqrt{(1+\delta)n\log n}}{(1-\delta)\sqrt{M}}\bigg{)}.$ (49)
Appendix E Proof of Theorems 3 and 4
Proof.
We begin by proving Theorem 4, after which we prove Theorem 3. For any density matrices $\bm{\rho}_{1},\bm{\rho}_{2}\in\mathbb{X}_{\bm{r}}$ in (11), we have
$\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}$ $=$ $\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}_{1}-\bm{\rho}_{2}\rangle$ (53)
$=$ $\max_{\mbox{\tiny$\begin{array}[]{c}\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}},\bm{\rho}=\bm{\rho}^{\dagger},\\ \text{rank}(\bm{\rho})=(2r_{1},\dots,2r_{n-1}),\\ \operatorname{trace}(\bm{\rho})=0,\|\bm{\rho}\|_{F}\leq\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}\end{array}$}}\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}\rangle.$
Furthermore, we define a restricted Frobenius norm as following:
$\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F,2\overline{r}}=\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}\rangle.$ (54)
Now, we can expand $\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$ as following:
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}=\|\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)}))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$ (55)
$\leq$ $(1+b)\|\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$
$=$ $(1+b)\bigg{(}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}-2\mu\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\nabla g(\bm{\rho}^{(\tau)})\rangle+\mu^{2}\|\nabla g(\bm{\rho}^{(\tau)})\|_{F,2\overline{r}}^{2}\bigg{)}.$
First, we need to analyze
$\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\nabla g(\bm{\rho}^{(\tau)})\rangle$ (56)
$=$ $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle^{2}-\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle$
$\geq$ $\frac{d^{n}(1-\delta)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}}{K(d^{n}+1)}-\frac{c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F},$
where the first inequality uses the Lemma 2 and (44) with probability $1-e^{-c_{2}nd^{2}\overline{r}^{2}\log n}$. Here $c_{1},c_{2}$ are positive constants.
Then we need to derive
$\|\nabla g(\bm{\rho}^{(\tau)})\|_{F,\overline{2}r}^{2}$$\leq$ $2\bigg{\|}\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle\bm{A}_{k}\bigg{\|}_{F,2\overline{r}}^{2}+2\bigg{\|}\sum_{k=1}^{K}\eta_{k}\bm{A}_{k}\bigg{\|}_{F,2\overline{r}}^{2}$ (57)
$\leq$ $\frac{d^{2n+1}(1+\delta)^{2}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}}{K^{2}(d^{n}+1)^{2}}+\frac{2c_{1}^{2}(1+\delta)\overline{r}^{2}nd^{n+2}\gamma_{t}(\bm{\rho}^{\star})\log n}{K^{2}(d^{n}+1)M},$
where the second inequality respectively follows (44) and
$\bigg{\|}\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle\bm{A}_{k}\bigg{\|}_{F,2\overline{r}}$ $=$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}\bigg{\langle}\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle\bm{A}_{k},\widetilde{\bm{\rho}}\bigg{\rangle}$ (58)
$=$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle\langle\bm{A}_{k},\widetilde{\bm{\rho}}\rangle$
$\leq$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}\frac{d^{n}(1+\delta)(\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle+\operatorname{trace}(\bm{\rho}^{(\tau)}-\bm{\rho}^{\star})\operatorname{trace}(\widetilde{\bm{\rho}}))}{K(d^{n}+1)}$
$\leq$ $\frac{d^{n}(1+\delta)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}}{K(d^{n}+1)},$
where the last line uses Lemma 8.
Combining (56) and (57), we arrive at
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$ (59)
$\leq$ $(1+b)\big{(}\big{(}1-\frac{d^{n+1}(1-\delta)\mu}{K(d^{n}+1)}+\frac{d^{2n+1}(1+\delta)^{2}\mu^{2}}{K^{2}(d^{n}+1)^{2}}\big{)}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big{)}$
$\leq$ $(1+b)\big{(}\big{(}1-\frac{d^{n}(1-\delta)\mu}{K(d^{n}+1)}\big{)}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big{)}$
$\leq$ $a^{\tau+1}\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}+\frac{1+b}{1-a}c_{3}\bigg{(}\frac{\mu d\overline{r}\sqrt{n(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{\mu^{2}(1+\delta)\overline{r}^{2}nd^{2}\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$
$\leq$ $a^{\tau+1}\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}+c_{4}\bigg{(}\frac{d\overline{r}\sqrt{n(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{(1+\delta)\overline{r}^{2}nd^{2}\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)},$
where we define the error term $e=\frac{2\mu c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{2\mu^{2}c_{1}^{2}(1+\delta)\overline{r}^{2}nd^{n+2}\gamma_{t}(\bm{\rho}^{\star})\log n}{K^{2}(d^{n}+1)M}$, and $c_{3},c_{4}$ are positive constants. The last two lines follow assumptions with regard to the step size and initialization: $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ and $b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}$ which guarantee that the convergence rate satisfies
$a=(1+b)\bigg{(}1-\frac{d^{n}(1-\delta)\mu}{K(d^{n}+1)}\bigg{)}<1.$ (60)
According to (59), the linear convergence of the PGD algorithm is formally established. Furthermore, when the number of iterations is sufficiently large such that $a^{\tau+1}\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}\leq c_{4}\frac{(1+\delta)\overline{r}^{2}nd^{2}\gamma_{t}(\bm{\rho}^{\star})\log n}{M}$, the recovery error of the PGD algorithm satisfies:
$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}\leq c_{5}\sqrt{\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}},$ (61)
where $c_{5}$ is a positive constant. This completes the proof of Theorem 4.
Next, we begin by proving Theorem 3. According to Lemma 5, we have
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}=\|\mathcal{P}_{\operatorname{trace}}(\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$ (62)
$\leq$ $\|\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)}))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$
$\leq$ $\bigg{(}1+\frac{600n}{\underline{\sigma}({\bm{\rho}^{\star}})}\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}\bigg{)}\|\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}.$
As long as $\frac{600n}{\underline{\sigma}({\bm{\rho}^{\star}})}\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}\leq b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}$, this completes the proof of Theorem 3.
Appendix F Proof of Theorem 5
Proof.
By the definition of the restricted Frobenius norm (54), we can analyze
$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}$ $=$ $\bigg{\|}\mathcal{P}_{\operatorname{trace}}\bigg{(}\text{SVD}_{\bm{r}}^{tt}\bigg{(}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg{)}-\bm{\rho}^{\star}\bigg{)}\bigg{\|}_{F,2\overline{r}}$ (63)
$\leq$ $\bigg{\|}\text{SVD}_{\bm{r}}^{tt}\bigg{(}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg{)}-\bm{\rho}^{\star}\bigg{\|}_{F,2\overline{r}}$
$\leq$ $(1+\sqrt{n-1})\bigg{\|}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}-\bm{\rho}^{\star}\bigg{\|}_{F,2\overline{r}}$
$\leq$ $(1+\sqrt{n-1})\bigg{(}\frac{K(d^{n}+1)}{d^{n}}\bigg{\|}\sum_{k=1}^{K}\eta_{k}\bm{A}_{k}\bigg{\|}_{F,2\overline{r}}+\bigg{\|}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle\bm{A}_{k}-\bm{\rho}^{\star}\bigg{\|}_{F,2\overline{r}}\bigg{)}$
$\leq$ $(1+\sqrt{n-1})\bigg{(}\frac{c_{1}d\overline{r}\sqrt{(1+\delta)(d^{n}+1)n\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{d^{n}M}}+\delta\|\bm{\rho}^{\star}\|_{F}\bigg{)},$
where the first two inequalities respectively follows the nonexpansiveness property of the projection onto the convex set and the quasi-optimality property of TT-SVD projection [27]. In the last line, the upper bound of $\|\sum_{k=1}^{K}\eta_{k}\bm{A}_{k}\|_{F,2\overline{r}}$ can be directly obtained from (41) with probability $1-e^{-c_{2}nd^{2}\overline{r}^{2}\log n}$ with constants $c_{i},i=1,2$. Additionally, based on Lemma 8, we have
$\bigg{\|}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle\bm{A}_{k}-\bm{\rho}^{\star}\bigg{\|}_{F,2\overline{r}}$ (64)
$=$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}\bigg{\langle}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle\bm{A}_{k}-\bm{\rho}^{\star},\widetilde{\bm{\rho}}\bigg{\rangle}$
$=$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}(1+\delta)(\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle+\operatorname{trace}(\bm{\rho}^{\star})\operatorname{trace}(\widetilde{\bm{\rho}}))-\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle\leq\delta\|\bm{\rho}^{\star}\|_{F}.$
Appendix G Auxiliary Materials
Lemma 6.
([40, Lemma 3]) The integral over the Haar measure in (5) is given by
[…]
where $P_{\text{Sym}}$ is the projector over the symmetric subspace.
Lemma 7.
([33, Lemma 10]) For any ${\bm{A}}_{i},{\bm{A}}^{\star}_{i}\in\mathbb{R}^{r_{i-1}\times r_{i}},i=1,\dots,N$, we have
${\bm{A}}_{1}{\bm{A}}_{2}\cdots{\bm{A}}_{N}-{\bm{A}}_{1}^{\star}{\bm{A}}_{2}^{\star}\cdots{\bm{A}}_{N}^{\star}=\sum_{i=1}^{N}\bm{A}_{1}^{\star}\cdots\bm{A}_{i-1}^{\star}(\bm{A}_{i}-\bm{A}_{i}^{\star})\bm{A}_{i+1}\cdots\bm{A}_{N}.$ (66)
Lemma 8.
For arbitrary Hermitian matrix $\bm{\rho}_{1},\bm{\rho}_{2}\in\mathbb{C}^{d^{n}\times d^{n}}$ and a set of $\delta$-approximate $2$-design POVMs $\bm{A}_{k}=\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger},k\in[K]$ in Definition 3, we have
$\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}_{1}\rangle\langle\bm{A}_{k},\bm{\rho}_{2}\rangle\leq(1+\delta)\frac{d^{n}(\operatorname{trace}(\bm{\rho}_{1})\operatorname{trace}(\bm{\rho}_{2})+\operatorname{trace}(\bm{\rho}_{1}\bm{\rho}_{2}))}{K(d^{n}+1)}.$ (67)
Proof.
$\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}_{1}\rangle\langle\bm{A}_{k},\bm{\rho}_{2}\rangle$ $=$ $\sum_{k=1}^{K}\operatorname{trace}((\bm{A}_{k}\bm{\rho}_{1})\otimes(\bm{A}_{k}\bm{\rho}_{2}))$ (68)
$=$ $\sum_{k=1}^{K}\operatorname{trace}(\bm{A}_{k}^{\otimes 2}(\bm{\rho}_{1}\otimes\bm{\rho}_{2}))$
$\leq$ $(1+\delta)\frac{d^{n}\operatorname{trace}(({\bf I}+\bm{S})(\bm{\rho}_{1}\otimes\bm{\rho}_{2}))}{K(d^{n}+1)}$
$=$ $(1+\delta)\frac{d^{n}(\operatorname{trace}(\bm{\rho}_{1})\operatorname{trace}(\bm{\rho}_{2})+\operatorname{trace}(\bm{\rho}_{1}\bm{\rho}_{2}))}{K(d^{n}+1)},$
where the penultimate line follows (6) and (35), and the last line is from [15, Lemma 17]. ∎
As a direct result of [103, Theorem 1.3], we have
Lemma 9.
Let $n$ and $d$ be positive integers and $f:\mathbb{C}^{d^{n}}\to\mathbb{R}$ be a convex and Lipschitz function. Consider a vector $\bm{x}=\bm{x}_{1}\otimes\cdots\otimes\bm{x}_{n}\in\mathbb{C}^{d^{n}}$ where all $\bm{x}_{k}\in\mathbb{C}^{d}$ are independent random vectors on the unit sphere. Then, for every $0\leq t\leq 2(\operatorname{\mathbb{E}}|f(\bm{x})|^{2})^{\frac{1}{2}}$, we have
$\operatorname{\mathbb{P}}\left(|f(\bm{x})-\operatorname{\mathbb{E}}f(\bm{x})|\geq t\right)\leq 2e^{-\frac{(d-1)t^{2}}{nL^{2}}},$ (69)
where $L$ is the Lipschitz constant of $f(\bm{x})$.

### v4

@@
-thanks: corresponding author
+Sample-Efficient Quantum State Tomography for Structured Quantum States in One Dimension
 Abstract
@@
 Moreover, we propose a projected gradient descent (PGD) algorithm to solve the constrained least-squares problem and show that it can efficiently find an estimate with bounded recovery error when appropriately initialized.
-1 Introduction
+I Introduction
 The state of a quantum system composed of $n$ qudits (which are $d$-level quantum systems; qubits have $d=2$) is fully described by a density matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$ with unit trace and is positive semidefinite (PSD) [1].
@@
 The density matrix elements of such states can be expressed as the following matrix product [25, 26]
-$\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},$ (1)
+$\hskip-14.22636pt\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},$ (1)
 where $i_{k}$ ($k=1,2,\cdots n$) labels a basis of the $k$-th qubit, and $\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}}$ with $r_{0}=r_{n}=1$.
@@
 In addition, this work [33] provides a projected gradient descent (i.e., iterative hard thresholding) algorithm for performing the QST in practice, but does not provide a formal guarantee for that algorithm.
-1.1 Main contributions
+I-A Main contributions
 Our work aims to address two fundamental questions regarding the QST of states represented by MPOs:
@@
     If so, can we develop a provably correct algorithm to reconstruct MPOs from the obtained measurements?
-Table 1: Comparison with existing work for estimating $n$-qudits MPOs (with bond dimension $\overline{r}$) to achieve $\epsilon$-closeness in the Frobenius norm.
+TABLE I: Comparison with existing work for estimating $n$-qudits MPOs (with bond dimension $\overline{r}$) to achieve $\epsilon$-closeness in the Frobenius norm.
 The result for SIC/2-designs depends on $\gamma(\bm{\rho})=K\cdot\max_{k}p_{k}$, i.e., the probability distribution of measuring the state $\bm{\rho}$ with the corresponding POVM.
-POVM ##\## POVMs ##\## total state copies
+POVM #\# POVMs #\# total state copies
 Haar random ([33, Theorem 5]) $\Omega(nd^{2}\overline{r}^{2}\log n)$ $O(\frac{n^{3}d^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
-SIC/$2$-designs (Theorem 1)             $1$ $O(\frac{nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n}{\epsilon^{2}})$
-$t$-designs, $t\geq 3$ (Theorem 2)             $1$ $O(\frac{nd^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
+SIC/22-designs (Theorem 1)        11 $O(\frac{nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n}{\epsilon^{2}})$
+$t$-designs, $t\geq 3$ (Theorem 2)        11 $O(\frac{nd^{2}\overline{r}^{2}\log n}{\epsilon^{2}})$
 We provide positive answers to both the above questions by utilizing informationally complete POVMs (IC-POVM).
@@
 We begin by studying recovery guarantees for MPOs from empirical quantum measurements (physical measurements containing statistical errors) obtained using a class of IC-PVOMs including the rank-one symmetric IC-POVM (SIC-POVM) [38] and more broadly spherical $t$-design POVMs that are induced by (approximate) spherical $t$-designs [39, 40].
-The first contribution of this paper—presented in Section 3—is a new sample complexity bound (refers to the number of state copies) that matches the number of parameters in MPOs to guarantee bounded recovery error for a particular estimator—the solution to a constrained least-squares optimization problem.
+The first contribution of this paper—presented in Section III—is a new sample complexity bound (refers to the number of state copies) that matches the number of parameters in MPOs to guarantee bounded recovery error for a particular estimator—the solution to a constrained least-squares optimization problem.
 Specifically, for SIC-POVM and $2$-design POVMs, our result shows that $M=O(nd^{2}\gamma(\bm{\rho})\overline{r}^{2}\log n/\epsilon^{2})$ number of state copies is sufficient to achieve $\epsilon$-accuracy in recovering the ground-truth state $\bm{\rho}$ in the Frobenius norm, where $\gamma(\bm{\rho})=K\cdot\max_{k=1}^{K}p_{k}$ captures the uniformity of the probabilistic distribution of the measurement.
@@
 This result improves upon previous work [33] mentioned above using Haar random projective measurement bases that requires $\Omega(nd^{2}\overline{r}^{2}\log n)$ POVMs and $O(n^{3}d^{2}\overline{r}^{2}\log n/\epsilon^{2})$ total number of state copies.
-We summarize the main results and comparison with our previous result [33] in Table 1.
+We summarize the main results and comparison with our previous result [33] in Table I.
 Moreover, physical MPO states are often approximately low-rank, allowing the previous analysis to be extended to the trace distance.
 Specifically, to ensure a bounded recovery error in terms of trace distance for rank-$s$ MPO states, the required number of state copies remains proportional to the degrees of freedom, $O(nsd^{2}\overline{r}^{2})$, when using (approximate) spherical $t$-designs with $t\geq 2$.
-The second contribution of this paper—presented in Section 4—is the development of a projected gradient descent (PGD) with guaranteed convergence analysis for solving the aforementioned constrained least-squares optimization problem.
+The second contribution of this paper—presented in Section IV—is the development of a projected gradient descent (PGD) with guaranteed convergence analysis for solving the aforementioned constrained least-squares optimization problem.
 In each iteration, the algorithm performs a gradient update, followed by a projection back to the MPO format with a unit trace.
@@
 While TT-SVD is not optimal, it is computationally efficient and comes with a sub-optimal guarantee.
-Leveraging this guarantee, we show that, with an appropriate initialization, PGD converges linearly to a vicinity of the ground-truth state, with a distance matching the results in Table 1.
+Leveraging this guarantee, we show that, with an appropriate initialization, PGD converges linearly to a vicinity of the ground-truth state, with a distance matching the results in Table I.
 In addition, we show that spectral initialization provides a valid starting point, ensuring the convergence of PGD.
-Our numerical experiments in Section 5 suggest that PGD can also efficiently find the underlying state even with a random initialization.
+Our numerical experiments in Section V suggest that PGD can also efficiently find the underlying state even with a random initialization.
 In this paper, we primarily focus on global spherical $t$-design POVMs with $t\geq 2$, as we theoretically demonstrate that stable recovery of constrained least-squares optimization for all MPO states can be achieved using a polynomial number of state copies.
@@
 In practice, the SIC-POVM or the spherical $t$-designs with $t\geq 2$ have no known efficient implementation using local quantum circuits.
-To partially address this practicality concern, we provide numerical evidence that local SIC-POVMs, which can be implemented efficiently and readily on current quantum hardware, can be used to perform QST for typical MPO states with $O(n)$ state copies using the proposed PGD algorithm in Section 5.
-1.2 Notation
+To partially address this practicality concern, we provide numerical evidence that local SIC-POVMs, which can be implemented efficiently and readily on current quantum hardware, can be used to perform QST for typical MPO states with $O(n)$ state copies using the proposed PGD algorithm in Section V.
+I-B Notation
 We use calligraphic letters (e.g., $\mathcal{X}$) to denote tensors, bold capital letters (e.g., $\bm{X}$) to denote matrices, bold lowercase letters (e.g., $\bm{x}$) to denote column vectors, and italic letters (e.g., $x$) to denote scalar quantities.
@@
 $\sigma_{i}(\bm{X})$ is the $i$-th singular value of $\bm{X}$.
-For a vector $\bm{a}$ of size $N\times 1$, its $l_{n}$-norm is defined as $||\bm{a}||_{n}=(\sum_{m=1}^{N}|a_{m}|^{n})^{\frac{1}{n}}$.
+For a vector $\bm{a}$ of size $N\times 1$, its $l_{n}$-norm is defined as $\bm{a}||_{n}=(\sum_{m=1}^{N}|a_{m}|^{n})^{\frac{1}{n}}$.
 For two positive quantities $a,b\in\mathbb{R}$, the inequality $b\lesssim a$ or $b=O(a)$ means $b\leq ca$ for some universal constant $c$; likewise, $b\gtrsim a$ or $b=\Omega(a)$ represents $b\geq ca$ for some universal constant $c$.
 To simplify notations in the following sections, for an order-$n$ TT format with ranks $(r_{1},\dots,r_{n})$, we define $\overline{r}=\max_{i=1}^{n-1}r_{i}$.
-2 Informationally Complete POVM
+II Informationally Complete POVM
 As described in the introduction, the population measurements $\{p_{k}\}$ of the density matrix $\bm{\rho}$ from a POVM $\{\bm{A}_{k}\}$ are linear measurements that can be described through a linear map $\mathcal{A}:\mathbb{C}^{d^{n}\times d^{n}}\rightarrow\mathbb{R}^{K}$ of the form
@@
 Suppose that $\{\bm{A}_{k}\}_{k=1}^{d^{2n}}$ form an SIC-POVM. Then for any Hermitian matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$, $\mathcal{A}(\bm{\rho})$ satisfies
-$\|\mathcal{A}(\bm{\rho})\|_{2}^{2}=\sum_{k=1}^{d^{2n}}\langle\bm{A}_{k},\bm{\rho}\rangle^{2}=\frac{\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}(\bm{\rho}))^{2}}{d^{n}(d^{n}+1)}.$ (3)
+$\hskip-17.07182pt\|\mathcal{A}(\bm{\rho})\|_{2}^{2}=\sum_{k=1}^{d^{2n}}\langle\bm{A}_{k},\bm{\rho}\rangle^{2}=\frac{\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}(\bm{\rho}))^{2}}{d^{n}(d^{n}+1)}.$ (3)
 The proof is provided in Appendix A.
@@
 In contrast, with POVM being informationally complete, (3) holds for any Hermitian matrices without relying on any randomness in the measurement operator.
-This allows us to develop tight sample complexity bound in Section 3 and iterative algorithms with guaranteed convergence in Section 4.
+This allows us to develop tight sample complexity bound in Section III and iterative algorithms with guaranteed convergence in Section IV.
 As an immediate consequence of Lemma 1, for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, the difference $\bm{\rho}_{1}-\bm{\rho}_{2}$ remains Hermitian, and hence
@@
 Definition 2.
-(Spherical $t$-designs [62, 15]).
-A finite set $\{\bm{w}_{k}\}_{k=1}^{K}\subset\mathbb{C}^{d^{n}}$ of normalized vectors is called a spherical quantum $t$-design if 444 $K\geq C_{d^{n}+\lfloor t/2\rfloor-1}^{\lfloor t/2\rfloor}\cdot C_{d^{n}+\lceil t/2\rceil-1}^{\lceil t/2\rceil}$ is necessary to form a spherical $t$-design[52, 63].
 […]
+$\frac{1}{K}\sum_{k=1}^{K}(\bm{w}_{k}\bm{w}_{k}^{\dagger})^{\otimes s}=\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w}$ (5)
 holds for any $s\leq t$, where the integral on the right hand side is taken with respect to the Haar measure on the complex unit sphere in $\mathbb{C}^{d^{n}}$.
-[…]
+When $s=1$, we have $\frac{1}{K}\sum_{k=1}^{K}\bm{w}_{k}\bm{w}_{k}^{\dagger}=\int\bm{w}\bm{w}^{\dagger}d\bm{w}=\frac{1}{d^{n}}{\bf I}$, and hence $\bm{A}_{k}=\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger},k=1,\dots,K$ form a rank-one POVM.
+For simplicity, we call such an induced POVM $\{\bm{A}_{k}=\frac{d^{n}}{K}\bm{w}_{k}\bm{w}_{k}^{\dagger}\}$ as a $t$-design POVM.
+SIC-POVM is a special instance of $2$-design POVMs.
+In (5), we adopt the setting of uniform weights ($1/K$ for each $\bm{w}_{k}$, which is mostly commonly used in practice) to simplify the analysis, but one can consider a more general scenario of varied weights [40].
+Unlike SIC-POVM, $t$-designs always exist and can, in principle, be constructed in any dimension and for any $t$ [64, 65], although in some cases these constructions can be inefficient, as they require vector sets of exponential size [66].
+Instead, we may use approximate $t$-designs which can be efficiently implemented in practice [67].
 Definition 3.
@@
 A finite set $\{\bm{w}_{k}\}_{k=1}^{K}\subset\mathbb{C}^{d^{n}}$ of normalized vectors is called a spherical quantum $\delta$-approximate $t$-design if
-[…]
-[…]
+$(1-\delta)\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w}\leq\frac{1}{K}\sum_{k=1}^{K}(\bm{w}_{k}\bm{w}_{k}^{\dagger})^{\otimes s}\leq(1+\delta)\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w}$ (6)
 holds for any $s\geq 2$, where the integral on the right hand side is taken with respect to the Haar measure on the complex projective space.
@@
 Then for arbitrary Hermitian matrix $\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}$, $\mathcal{A}(\bm{\rho})$ satisfies
-$(1-\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}\leq\|\mathcal{A}(\bm{\rho})\|_{2}^{2}$
-$\leq(1+\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}.$ (7)
-The proof is given in Appendix B. Similar to (4), it follows from Lemma 2 that for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, $\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})$ satisfies
+$(1-\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}\leq\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\leq(1+\delta)\frac{d^{n}(\|\bm{\rho}\|_{F}^{2}+(\operatorname{trace}({\bm{\rho}}))^{2})}{K(d^{n}+1)}.$ (7)
+The proof is given in Appendix B. Similar to (4), it follows from Equation 7 that for any two density matrices $\bm{\rho}_{1}$ and $\bm{\rho}_{2}$, $\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})$ satisfies
 $\|\mathcal{A}(\bm{\rho}_{1}-\bm{\rho}_{2})\|_{2}^{2}$ (8)
@@
 $=$ $(1-\delta)\frac{d^{n}\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}}{K(d^{n}+1)}.$
-3 Stable Recovery of Matrix Product Operators
+III Stable Recovery of Matrix Product Operators
 Building upon the results that a (approximate) $t$-design POVM $\{\bm{A}_{1},\dots,\bm{A}_{K}\}\in\mathbb{C}^{d^{n}\times d^{n}}$ provides unique population measurements for any state, in this section, we investigate the stable recovery for MPOs of form (1) from empirical measurements acquired through such POVMs.
@@
 With empirical measurements $\widehat{\bm{p}}$, we consider minimizing the following constrained least squares objective:
-$\widehat{\bm{\rho}}=\operatorname*{\text{arg\leavevmode\nobreak\ min}}_{\bm{\rho}\in\mathbb{X}_{\bm{r}}}\|\mathcal{A}(\bm{\rho})-{\widehat{\bm{p}}}\|_{2}^{2},$ (10)
+$\widehat{\bm{\rho}}=\operatorname*{\text{arg~min}}_{\bm{\rho}\in\mathbb{X}_{\bm{r}}}\|\mathcal{A}(\bm{\rho})-{\widehat{\bm{p}}}\|_{2}^{2},$ (10)
 where we denote by $\mathbb{X}_{\bm{r}}$ the set of MPOs with MPO ranks $\bm{r}=(r_{1},\dots,r_{n-1})$:
@@
 Using the fact that $\widehat{\bm{\rho}}$ is a global solution of (10) and the fact that $\bm{\rho}^{\star}\in\mathbb{X}_{\bar{r}}$, we have
-00 0 $\leq$ $\|\mathcal{A}(\bm{\rho}^{\star})-\widehat{\bm{p}}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\widehat{\bm{p}}\|_{2}^{2}$ (12)
+$0$ $\!\!\!\!\leq\!\!\!\!$ $\|\mathcal{A}(\bm{\rho}^{\star})-\widehat{\bm{p}}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\widehat{\bm{p}}\|_{2}^{2}$ (12)
 $\!\!\!\!=\!\!\!\!$ $\|\mathcal{A}(\bm{\rho}^{\star})-\mathcal{A}(\bm{\rho}^{\star})-\bm{\eta}\|_{2}^{2}-\|\mathcal{A}(\widehat{\bm{\rho}})-\mathcal{A}(\bm{\rho}^{\star})-\bm{\eta}\|_{2}^{2}$
@@
 Then, for any $t>0$,
-$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}a_{k}(\frac{f_{k}}{M}-p_{k})>t\right)$
-$\leq e^{-\frac{Mt}{4a_{\max}}\min\bigg{\{}1,\frac{a_{\max}t}{4\sum_{k=1}^{K}a_{k}^{2}p_{k}}\bigg{\}}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}a_{k}^{2}p_{k}}},$
+$\operatorname{\mathbb{P}}\left(\sum_{k=1}^{K}a_{k}(\frac{f_{k}}{M}-p_{k})>t\right)\leq e^{-\frac{Mt}{4a_{\max}}\min\bigg\{1,\frac{a_{\max}t}{4\sum_{k=1}^{K}a_{k}^{2}p_{k}}\bigg\}}+e^{-\frac{Mt^{2}}{8\sum_{k=1}^{K}a_{k}^{2}p_{k}}},$ (15)
 where $a_{\max}=\max_{k}|a_{k}|$.
 We proceed by considering the following two cases, $2$-designs and $t$-designs with $t\geq 3$.
-Case $(i)$: $2$-designs.
-There are two potential technical difficulties in directly applying Lemma 3 to bound the term $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ by plugging in $a_{k}=\langle\bm{A}_{k},\widehat{\bm{\rho}}-\bm{\rho}^{\star}\rangle$: (1) $a_{k}$ in not independent to the random variables $\{f_{k}\}$ since $\widehat{\bm{\rho}}$ depends on the empirical measurements, and (2) we need to upper bound a term of form $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ originating from the exponential term in (3).
+Case $(i)$: 22-designs.
+There are two potential technical difficulties in directly applying Equation 15 to bound the term $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ by plugging in $a_{k}=\langle\bm{A}_{k},\widehat{\bm{\rho}}-\bm{\rho}^{\star}\rangle$: (1) $a_{k}$ in not independent to the random variables $\{f_{k}\}$ since $\widehat{\bm{\rho}}$ depends on the empirical measurements, and (2) we need to upper bound a term of form $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}p_{k}$ originating from the exponential term in (15).
 We will address the first challenge by using a covering argument.
-For the second challenge, to utlize the upper bound for $\|\mathcal{A}(\bm{\rho}-\bm{\rho}^{\star})\|_{2}^{2}=\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}$ in Section 2, we capture the largest probabilities by
+For the second challenge, to utilize the upper bound for $\|\mathcal{A}(\bm{\rho}-\bm{\rho}^{\star})\|_{2}^{2}=\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}-\bm{\rho}^{\star}\rangle^{2}$ in Section II, we capture the largest probabilities by
 $\gamma(\bm{\rho}^{\star}):=K\cdot\max_{k}p_{k}=K\cdot\max_{k}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle.$ (16)
@@
 However, this work [74] demonstrates one method to generate spherical 3-designs by extracting them from unitary designs.
-On the other hand, the SIC-POVM is not a $3$-design POVM, hence the assumption of "uniformly" distributed probabilities may be indispensable.
+On the other hand, the SIC-POVM is not a $3$-design POVM, hence the assumption of ”uniformly” distributed probabilities may be indispensable.
 While the solution $\widehat{\bm{\rho}}$ of (10) may be non-physical, we can impose additional PSD constraint to obtain a physical state without compromising the recovery guarantee in Theorem 1.
@@
 Since the set $\mathbb{S}_{+}$ is convex, the corresponding projector is non-expansive, and hence
-$\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-\bm{\rho}^{\star}\|_{F}$ $=$ $\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-P_{\mathbb{S}_{+}}(\bm{\rho}^{\star})\|_{F}$ (19)
-$\leq$ $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq\epsilon,$
+$\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-\bm{\rho}^{\star}\|_{F}=\|P_{\mathbb{S}_{+}}(\widehat{\bm{\rho}})-P_{\mathbb{S}_{+}}(\bm{\rho}^{\star})\|_{F}\leq\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\leq\epsilon,$ (19)
 which implies that the projection step ensures that the state becomes physically valid while preserving or even improving the recovery guarantee.
@@
 If we measure the state $M$ times using the POVM, then with high probability, the solution $\widehat{\bm{\rho}}$ of the constrained least squares (10) satisfies
-$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}$ $\lesssim\frac{d\overline{r}\sqrt{(1+\delta)ns\gamma_{t}(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}}$ (20)
-$+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1}.$
+$\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{1}\lesssim\frac{d\overline{r}\sqrt{(1+\delta)ns\gamma_{t}(\bm{\rho}^{\star})\log n}}{(1-\delta)\sqrt{M}}+2\sqrt{s}\|\bm{N}\|_{F}+\|\bm{N}\|_{1}.$ (20)
 As shown in Corollary 1, when the ground truth MPO state is approximately low-rank, the recovery error in terms of trace distance also exhibits a polynomial growth with respect to $n$.
@@
 $\inf_{\widehat{\bm{\rho}}}\sup_{\bm{\rho}^{\star}\in\mathbb{X}_{\bm{r}}}\operatorname{\mathbb{E}}{\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}}\geq\Omega\bigg(\sqrt{\frac{nd^{2}\overline{r}^{2}}{M}}\bigg).$ (21)
-Compared to (21), the recovery upper bound in Theorem 2 is tight and optimal, up to logarithmic terms.
+Compared to (21), the recovery upper bound in Theorem 2 is tight, up to logarithmic terms.
 We now briefly discuss technical challenges in extending the above minimax lower bound to the statistical noise following multinomial distribution.
@@
 Moreover, recovering an MPO state is expected to require much more state copies than testing whether an unknown state is an MPS.
-4 Projected Gradient Descent with Guaranteed Convergence
+IV Projected Gradient Descent with Guaranteed Convergence
 In the literature of tensor recovery for tensor train (TT) format tensors, several iterative algorithms [83, 84, 85, 79, 86] have been proposed and analyzed from random measurements.
@@
 Lemma 5.
-([88, Lemma 26]) Let $\bm{\rho}^{\star}\in\mathbb{C}^{d^{n}\times d^{n}}$ be an MPO with MPO ranks $(r_{1},\dots,r_{n-1})$.
-Denote by $\underline{\sigma}({\bm{\rho}^{\star}})$ the smallest TT singular value of $\bm{\rho}^{\star}$ [^15] Specifically, we reshape $\bm{\rho}$ into an $n$-th order tensor $\mathcal{X}$ of size $d^{2}\times d^{2}\times\cdots\times d^{2}$ and define the $\ell$-th unfolding matrix of tensor $\mathcal{X}$ as $\bm{X}^{\langle\ell\rangle}\in\mathbb{C}^{d^{2\ell}\times d^{2n-2\ell}}$, where the $(s_{1}\cdots s_{\ell},s_{\ell+1}\cdots s_{n})$-th element of $\bm{X}^{\langle\ell\rangle}$ is given by $\bm{X}^{\langle\ell\rangle}(s_{1}\cdots s_{\ell},s_{\ell+1}\cdots s_{n})=\mathcal{X}(s_{1},\dots,s_{n})$.
-With the $\ell$-th unfolding matrix $\mathcal{X}^{\langle\ell\rangle}$ and the MPO ranks, we can obtain its smallest TT singular value $\underline{\sigma}(\mathcal{X})=\min_{\ell=1}^{n-1}\sigma_{r_{\ell}}(\mathcal{X}^{\langle\ell\rangle})$..
-Then for any perturbation $\bm{E}\in\mathbb{C}^{d^{n}\times d^{n}}$ with $C_{n}||\bm{E}||_{F}\leq\underline{\sigma}({\bm{\rho}^{\star}})$ for some constant $C_{n}\geq 500n$, we have
+[…]
 $\|\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{\star}+\bm{E})-\bm{\rho}^{\star}\|_{F}^{2}\leq\|\bm{E}||_{F}^{2}+\frac{600n||\bm{E}\|_{F}^{3}}{\underline{\sigma}({\bm{\rho}^{\star}})}.$
@@
 Given an initialization $\bm{\rho}^{(0)}$ satisfying
-$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}<\frac{\underline{\sigma}({\bm{\rho}^{\star}})(d-1)(1-\delta)^{2}}{600n(1+\delta^{2}+(4d-2)\delta)},$ (24)
+$\hskip-28.45274pt\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}<\frac{\underline{\sigma}({\bm{\rho}^{\star}})(d-1)(1-\delta)^{2}}{600n(1+\delta^{2}+(4d-2)\delta)},$ (24)
 then with probability at least $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, the PGD in (23) with step size $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ converges at a linear rate and
-$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg{(}\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$ (25)
+$\hskip-22.76228pt\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg(\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg)$ (25)
 for sufficiently large $\tau$.
@@
 However, this is not the case for many practical states, such as those at low temperatures or systems close to pure states.
-However, as will be illustrated in Section 5, our experiments show that even a randomly initialized PGD can efficiently converge to a stable recovery error.
+However, as will be illustrated in Section V, our experiments show that even a randomly initialized PGD can efficiently converge to a stable recovery error.
 To relax the requirement on the initialization, the following result establishes another convergence guarantee with an additional requirement on the projection.
@@
 Given an initialization $\bm{\rho}^{(0)}$, suppose the PGD in (23) uses step size $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ and that the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}$ satisfies
-$\|\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F}^{2}\leq(1+b)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2},$
-$\text{with}\ b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}.$ (26)
+$\|\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\bm{\rho}^{(\tau)})-\bm{\rho}^{\star}\|_{F}^{2}\leq(1+b)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2},\text{with}\ b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}.$ (26)
 Then with probability at least $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, the PGD converges at a linear rate and
-$\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg{(}\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg{)}$ (27)
+$\hskip-22.76228pt\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}\leq O\bigg(\frac{nd^{2}\overline{r}^{2}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg)$ (27)
 for sufficiently large $\tau$.
@@
 By utilizing the spectral initialization in (28), with probability $1-e^{-\Omega(nd^{2}\overline{r}^{2}\log n)}$, we have
-[…]
-$\cdot O\bigg{(}\frac{d\overline{r}\sqrt{(1+\delta)n\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}+\delta\|\bm{\rho}^{\star}\|_{F}\bigg{)}.$
+$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}\leq(1+\sqrt{n-1})\cdot O\bigg(\frac{d\overline{r}\sqrt{(1+\delta)n\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}+\delta\|\bm{\rho}^{\star}\|_{F}\bigg).$ (29)
 Its proof is given in Appendix F.
@@
 Note that the requirements on the number of measurements are slightly more stringent than those in Theorem 1 and Theorem 2 due to the sub-optimal bound for the projection $\mathcal{P}_{\mathbb{X}_{\bm{r}}}(\cdot)$.
-However, as will be illustrated in Section 5, our numerical experiments show that even a randomly initialized PGD can very efficiently find an estimate of the target state.
+However, as will be illustrated in Section V, our numerical experiments show that even a randomly initialized PGD can very efficiently find an estimate of the target state.
 A theoretical explanation for this phenomenon will be the focus of future research.
@@
 For instance, each time (often called epoch in machine learning), we first select a subset (say $N$) of empirical probabilities that include all the nonzero ones and a few numbers of zero ones (which can be either randomly selected or selected in a sequential), and then perform $N/B$ iterations of PSGD with each iteration using $B$ measurements selected sequentially form this subset.
-5 Numerical Experiments with Local IC-POVMs
+V Numerical Experiments with Local IC-POVMs
 To our best knowledge, there are no known efficient implementations of SIC-POVM or spherical $t$-designs with $t\geq 2$ for a many-qubit system using local quantum circuits available in current or near-future quantum hardware.
@@
 Specifically, we consider the following local SIC-POVM $\{\bm{B}_{i}\}_{i=1}^{4}$ for each qubit
-$\bigg{\{}\begin{bmatrix}\frac{1}{2}&0\\ 0&0\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}\\ \frac{\sqrt{2}}{6}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{2\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{2\pi}{3}}&\frac{1}{6}\end{bmatrix},$
-$\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{4\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{4\pi}{3}}&\frac{1}{6}\end{bmatrix}\bigg{\}}.$ (30)
+$\bigg\{\begin{bmatrix}\frac{1}{2}&0\\ 0&0\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}\\ \frac{\sqrt{2}}{6}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{2\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{2\pi}{3}}&\frac{1}{6}\end{bmatrix},\begin{bmatrix}\frac{1}{6}&\frac{\sqrt{2}}{6}e^{-i\frac{4\pi}{3}}\\ \frac{\sqrt{2}}{6}e^{i\frac{4\pi}{3}}&\frac{1}{6}\end{bmatrix}\bigg\}$.
+(30)
 We can then generate the set of IC-POVM for an $n$-qubit system as $\{\bm{A}_{i}\}_{i=1}^{4^{n}}=\{\bm{B}_{i_{1}}\otimes\cdots\otimes\bm{B}_{i_{n}}\}_{i_{1},\dots,i_{n}}$.
@@
 Figure 1: Illustration of (a) recovery error for different initialization methods, $\overline{r}$ and $n$ with $M=3000$; (b) convergence rate of PGD for different $n$, spectral initialization (dotted lines), and random initialization (solid lines) with $M=3000$ and $\overline{r}=4$; (c) recovery error with $n=5$ for different initialization methods, $\overline{r}$ and $M$; (d) $p_{\text{max}}$ for the MPDO.
-While our main results (Theorems 1 and 2) do not apply to such a measurement setting, we show that certain MPO states can be recovered using this IC-POVM using a number of state copies only linear in $n$ and the PGD and PSGD algorithms developed in Section 3.
+While our main results (Theorems 1 and 2) do not apply to such a measurement setting, we show that certain MPO states can be recovered using this IC-POVM using a number of state copies only linear in $n$ and the PGD and PSGD algorithms developed in Section III.
 A more systematic study on what types of MPO states can be recovered efficiently using the local IC-POVM can be found in [98], using a different optimization algorithm based on maximal likelihood estimation (MLE).
@@
 Each entry of ${\bm{A}_{\ell}^{\star}}^{i_{\ell},a_{\ell}}$ is i.i.d.
-randomly generated with both its real and imaginary parts drawn uniformly from the interval $[-1,1]$.
+randomly generated with both its real and imaginary parts drawn uniformly from the interval [1,1][-1,1].
 In the experiments, we set $K_{\ell}=10$ to ensure the state is sufficiently mixed; using $K_{\ell}=1$ would generate a pure state.
@@
 To further elucidate the computational complexity of the two methods, we summarize the computational cost for a single iteration of PGD and one epoch of PSGD under local measurements as follows:
-[…]
-[…]
-6 Conclusion
+$\text{PGD:}\ O\bigg(\underbrace{4^{n}n{\overline{r}}^{3}}_{\text{inner product}}+\underbrace{4^{n}{\overline{r}}^{2}}_{\text{TT-SVD}}+\underbrace{n{\overline{r}}^{2}}_{\text{unit trace}}\bigg),$
+$\text{PSGD:}\ O\bigg(\frac{N}{B}(\underbrace{Bn{\overline{r}}^{3}}_{\text{inner product}}+\underbrace{n({\overline{r}}+B)^{3}}_{\text{rounding}}+\underbrace{n{\overline{r}}^{2}}_{\text{unit trace}})\bigg).$
+VI Conclusion
 This paper focuses on the sample complexity bounds for recovering structured quantum states in one-dimension that are represented as matrix product operators (MPOs) with finite bond dimensions.
@@
 $\operatorname{trace}(\bm{\rho}^{2})$ $\!\!\!\!=\!\!\!\!$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\sum_{j\neq k}\langle\bm{A}_{j},\bm{\rho}\rangle\langle\bm{A}_{k},\bm{\rho}\rangle$ (34)
-$=$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\bigg{(}\bigg{(}\sum_{i=1}^{d^{2n}}\langle\bm{A}_{i},\bm{\rho}\rangle\bigg{)}^{2}-\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\bigg{)}$
+$\!\!\!\!=\!\!\!\!$ $(d^{n}(d^{n}+1)-1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-\big(\big(\sum_{i=1}^{d^{2n}}\langle\bm{A}_{i},\bm{\rho}\rangle\big)^{2}-\|\mathcal{A}(\bm{\rho})\|_{2}^{2}\big)$
 $\!\!\!\!=\!\!\!\!$ $d^{n}(d^{n}+1)\|\mathcal{A}(\bm{\rho})\|_{2}^{2}-(\operatorname{trace}(\bm{\rho}))^{2},$
@@
 According to [40, Theorem 1] and Lemma 6, we have
-[…]
+$\int(\bm{w}\bm{w}^{\dagger})^{\otimes 2}d\bm{w}=\frac{{\bf I}+\bm{S}}{d^{n}(d^{n}+1)},$ (35)
 where $\bm{S}$ is the swap operator. Taking the inner product between (35) and $\frac{d^{2n}}{K}\bm{\rho}^{\otimes 2}$ we get
-[…]
+$\bigg\langle\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w},\frac{d^{2n}}{K}\bm{\rho}^{\otimes 2}\bigg\rangle$ $\!\!\!\!=\!\!\!\!$ $\frac{d^{n}(\operatorname{trace}(\bm{\rho}^{\otimes 2})+\operatorname{trace}(\bm{S}\bm{\rho}^{\otimes 2}))}{K(d^{n}+1)}$ (36)
 $\!\!\!\!=\!\!\!\!$ $\frac{d^{n}((\operatorname{trace}(\bm{\rho}))^{2}+\operatorname{trace}(\bm{\rho}^{2}))}{K(d^{n}+1)}$
@@
 Now, we start by upper-bounding $\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$. Towards that goal, we first rewrite this term as
-$\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle=\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle\leq\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle,$ (37)
+$\langle\bm{\eta},\mathcal{A}(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ $\!\!\!\!=\!\!\!\!$ $\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},(\widehat{\bm{\rho}}-\bm{\rho}^{\star})\rangle$ (37)
+$\!\!\!\!\leq\!\!\!\!$ $\|\widehat{\bm{\rho}}-\bm{\rho}^{\star}\|_{F}\max_{\bm{\rho}\in\overline{\mathbb{X}}_{2\bm{r}}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle,$
 where we denote by $\overline{\mathbb{X}}_{\bm{r}}$ the normalized set of MPOs with $\bm{r}=[r_{1},\dots,r_{n-1}]$:
-$\begin{split}\overline{\mathbb{X}}_{\bm{r}}=\Big{\{}&\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\ \bm{\rho}=\bm{\rho}^{\dagger},\|\bm{\rho}\|_{F}\leq 1,\operatorname{trace}(\bm{\rho})=0,\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},\\ &\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}},\ell\in[n-1],r_{0}=r_{n}=1\Big{\}}.\end{split}$ (38)
+$\begin{split}\overline{\mathbb{X}}_{\bm{r}}=\Big\{&\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}}:\ \bm{\rho}=\bm{\rho}^{\dagger},\|\bm{\rho}\|_{F}\leq 1,\operatorname{trace}(\bm{\rho})=0,\\ &\bm{\rho}(i_{1}\cdots i_{n},j_{1}\cdots j_{n})=\bm{X}_{1}^{i_{1},j_{1}}\bm{X}_{2}^{i_{2},j_{2}}\cdots\bm{X}_{n}^{i_{n},j_{n}},\\ &\bm{X}_{\ell}^{i_{\ell},j_{\ell}}\in\mathbb{C}^{r_{\ell-1}\times r_{\ell}},\ell\in[n-1],r_{0}=r_{n}=1\Big\}.\end{split}$ (38)
 It is important to highlight that in comparison to $\mathbb{X}_{\bm{r}}$ in (11), $\overline{\mathbb{X}}_{\bm{r}}$ additionally includes $\|\bm{\rho}\|_{F}\leq 1$ and $\operatorname{trace}(\bm{\rho})=0$.
 The rest of the proof is to bound $\max_{\bm{\rho}}\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}\rangle$, which will be achieved by using a covering argument.
-First, when conditioned on $\bm{A}_{k}$, we consider any fixed value of $\bm{\rho}^{(p)}=[\bm{X}_{1}^{(p_{1})},\dots,\bm{X}_{n}^{(p_{n})}]\in\overline{\mathbb{X}}_{2\bm{r}}$ and according to [33, Appendix B-A], construct a $\xi$-net $\{\bm{\rho}^{(1)},\ldots,\bm{\rho}^{(N_{1}\cdots N_{n})}\}=\{[\bm{X}_{1}^{(1)},\dots,\bm{X}_{n}^{(1)}],\dots,[\bm{X}_{1}^{(N_{1})},\dots,\bm{X}_{n}^{(N_{n})}]\}$ such that $\sup_{\bm{X}_{\ell}:\|\bm{X}_{\ell}\|_{F}\leq 1}\leavevmode\nobreak\ \min_{p_{\ell}\leq N_{\ell}}\|\bm{X}_{\ell}-\bm{X}_{\ell}^{(p_{\ell})}\|_{F}\leq\xi,\ell=1,\dots,n$ with covering number $\Pi_{\ell=1}^{n}N_{\ell}\leq(\frac{4+\xi}{\xi})^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}$.
+First, when conditioned on $\bm{A}_{k}$, we consider any fixed value of $\bm{\rho}^{(p)}=[\bm{X}_{1}^{(p_{1})},\dots,\bm{X}_{n}^{(p_{n})}]\in\overline{\mathbb{X}}_{2\bm{r}}$ and according to [33, Appendix B-A], construct a $\xi$-net $\{\bm{\rho}^{(1)},\ldots,\bm{\rho}^{(N_{1}\cdots N_{n})}\}=\{[\bm{X}_{1}^{(1)},\dots,\bm{X}_{n}^{(1)}],\dots,[\bm{X}_{1}^{(N_{1})},\dots,\bm{X}_{n}^{(N_{n})}]\}$ such that $\sup_{\bm{X}_{\ell}:\|\bm{X}_{\ell}\|_{F}\leq 1}~\min_{p_{\ell}\leq N_{\ell}}\|\bm{X}_{\ell}-\bm{X}_{\ell}^{(p_{\ell})}\|_{F}\leq\xi,\ell=1,\dots,n$ with covering number $\Pi_{\ell=1}^{n}N_{\ell}\leq(\frac{4+\xi}{\xi})^{8r_{1}+\sum_{i=2}^{n-1}16r_{i-1}r_{i}+8r_{n-1}}$.
 Then we apply Lemma 3 to establish a concentration inequality for the expression $\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(p)}\rangle$ as follows:
@@
 $\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}^{2}$ $=$ $\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}_{1}-\bm{\rho}_{2}\rangle$ (53)
-$=$ $\max_{\mbox{\tiny$\begin{array}[]{c}\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}},\bm{\rho}=\bm{\rho}^{\dagger},\\ \text{rank}(\bm{\rho})=(2r_{1},\dots,2r_{n-1}),\\ \operatorname{trace}(\bm{\rho})=0,\|\bm{\rho}\|_{F}\leq\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}\end{array}$}}\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}\rangle.$
+$=$ $\hskip-42.67912pt\max_{\mbox{\tiny$\begin{array}[]{c}\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}},\bm{\rho}=\bm{\rho}^{\dagger},\\ \text{rank}(\bm{\rho})=(2r_{1},\dots,2r_{n-1}),\\ \operatorname{trace}(\bm{\rho})=0,\|\bm{\rho}\|_{F}\leq\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}\end{array}$}}\hskip-28.45274pt\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}\rangle.$
 Furthermore, we define a restricted Frobenius norm as following:
@@
 First, we need to analyze
-$\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\nabla g(\bm{\rho}^{(\tau)})\rangle$ (56)
-$=$ $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle^{2}-\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle$
+$\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\nabla g(\bm{\rho}^{(\tau)})\rangle$ $\!\!\!\!=\!\!\!\!$ $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle^{2}-\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle$ (56)
 $\!\!\!\!\geq\!\!\!\!$ $\frac{d^{n}(1-\delta)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}}{K(d^{n}+1)}-\frac{c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F},$
@@
 Combining (56) and (57), we arrive at
-$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$ (59)
-$\leq$ $(1+b)\big{(}\big{(}1-\frac{d^{n+1}(1-\delta)\mu}{K(d^{n}+1)}+\frac{d^{2n+1}(1+\delta)^{2}\mu^{2}}{K^{2}(d^{n}+1)^{2}}\big{)}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big{)}$
+$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$ $\!\!\!\!\leq\!\!\!\!$ $(1+b)\big(\big(1-\frac{d^{n+1}(1-\delta)\mu}{K(d^{n}+1)}+\frac{d^{2n+1}(1+\delta)^{2}\mu^{2}}{K^{2}(d^{n}+1)^{2}}\big)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big)$ (59)
 $\!\!\!\!\leq\!\!\!\!$ $(1+b)\big(\big(1-\frac{d^{n}(1-\delta)\mu}{K(d^{n}+1)}\big)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big)$
@@
 $\!\!\!\!\leq\!\!\!\!$ $a^{\tau+1}\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}^{2}+c_{4}\bigg(\frac{d\overline{r}\sqrt{n(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{\sqrt{M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{(1+\delta)\overline{r}^{2}nd^{2}\gamma_{t}(\bm{\rho}^{\star})\log n}{M}\bigg),$
-where we define the error term $e=\frac{2\mu c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{2\mu^{2}c_{1}^{2}(1+\delta)\overline{r}^{2}nd^{n+2}\gamma_{t}(\bm{\rho}^{\star})\log n}{K^{2}(d^{n}+1)M}$, and $c_{3},c_{4}$ are positive constants.
+where we define the error term
+$e=\frac{2\mu c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{2\mu^{2}c_{1}^{2}(1+\delta)\overline{r}^{2}nd^{n+2}\gamma_{t}(\bm{\rho}^{\star})\log n}{K^{2}(d^{n}+1)M},$ (60)
+and $c_{3},c_{4}$ are positive constants.
 The last two lines follow assumptions with regard to the step size and initialization: $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ and $b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}$ which guarantee that the convergence rate satisfies
@@
 Next, we begin by proving Theorem 3. According to Lemma 5, we have
-$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}=\|\mathcal{P}_{\operatorname{trace}}(\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$ (62)
+$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$
+$\!\!\!\!=\!\!\!\!$ $\|\mathcal{P}_{\operatorname{trace}}(\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$
 $\!\!\!\!\leq\!\!\!\!$ $\|\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)}))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$
@@
 By the definition of the restricted Frobenius norm (54), we can analyze
-$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}$ $=$ $\bigg{\|}\mathcal{P}_{\operatorname{trace}}\bigg{(}\text{SVD}_{\bm{r}}^{tt}\bigg{(}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg{)}-\bm{\rho}^{\star}\bigg{)}\bigg{\|}_{F,2\overline{r}}$ (63)
+$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}$
+$\!\!\!\!=\!\!\!\!$ $\bigg\|\mathcal{P}_{\operatorname{trace}}\bigg(\text{SVD}_{\bm{r}}^{tt}\bigg(\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg)-\bm{\rho}^{\star}\bigg)\bigg\|_{F,2\overline{r}}$
 $\!\!\!\!\leq\!\!\!\!$ $\bigg\|\text{SVD}_{\bm{r}}^{tt}\bigg(\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg)-\bm{\rho}^{\star}\bigg\|_{F,2\overline{r}}$
@@
 $\!\!\!\!=\!\!\!\!$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}\bigg\langle\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle\bm{A}_{k}-\bm{\rho}^{\star},\widetilde{\bm{\rho}}\bigg\rangle$
-$=$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}(1+\delta)(\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle+\operatorname{trace}(\bm{\rho}^{\star})\operatorname{trace}(\widetilde{\bm{\rho}}))-\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle\leq\delta\|\bm{\rho}^{\star}\|_{F}.$
+$\!\!\!\!=\!\!\!\!$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}(1+\delta)(\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle+\operatorname{trace}(\bm{\rho}^{\star})\operatorname{trace}(\widetilde{\bm{\rho}}))-\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle$
+$\!\!\!\!\leq\!\!\!\!$ $\delta\|\bm{\rho}^{\star}\|_{F}.$
 Appendix G Auxiliary Materials
@@
 ([40, Lemma 3]) The integral over the Haar measure in (5) is given by
-[…]
+$\int(\bm{w}\bm{w}^{\dagger})^{\otimes s}d\bm{w}=\frac{1}{C_{d^{n}+s-1}^{s}}P_{\text{Sym}},$ (66)
 where $P_{\text{Sym}}$ is the projector over the symmetric subspace.
@@
 ([33, Lemma 10]) For any ${\bm{A}}_{i},{\bm{A}}^{\star}_{i}\in\mathbb{R}^{r_{i-1}\times r_{i}},i=1,\dots,N$, we have
-${\bm{A}}_{1}{\bm{A}}_{2}\cdots{\bm{A}}_{N}-{\bm{A}}_{1}^{\star}{\bm{A}}_{2}^{\star}\cdots{\bm{A}}_{N}^{\star}=\sum_{i=1}^{N}\bm{A}_{1}^{\star}\cdots\bm{A}_{i-1}^{\star}(\bm{A}_{i}-\bm{A}_{i}^{\star})\bm{A}_{i+1}\cdots\bm{A}_{N}.$ (66)
+${\bm{A}}_{1}{\bm{A}}_{2}\cdots{\bm{A}}_{N}-{\bm{A}}_{1}^{\star}{\bm{A}}_{2}^{\star}\cdots{\bm{A}}_{N}^{\star}$ (67)
+$\!\!\!\!=\!\!\!\!$ $\sum_{i=1}^{N}\bm{A}_{1}^{\star}\cdots\bm{A}_{i-1}^{\star}(\bm{A}_{i}-\bm{A}_{i}^{\star})\bm{A}_{i+1}\cdots\bm{A}_{N}.$
 Lemma 8.